### v1

# The Discovery of Little Red Dots in the Local Universe: Signatures of Cool Gas Envelopes

Abstract
JWST observations have revealed a population of high-redshift “little red dots” (LRDs) that challenge conventional AGN models. We report the discovery of three local LRDs at –, initially selected from the SDSS database, with follow-up optical/near-IR spectroscopy and photometry. They exhibit properties fully consistent with those of high-redshift LRDs, including broad hydrogen and helium emission lines, compact morphologies, V-shaped UV-optical SED, declining near-IR continua, and no significant variability. Two sources were targeted but not detected in X-rays with statistical significance. All three sources show blue-shifted He I absorption, while two exhibit H and Na D absorption lines. We detect full Balmer and Paschen line series in all three objects, along with abundant narrow [Fe II] emission in two. The emission line analyses suggest narrow lines originate from AGN-powered, metal-poor regions with minimal dust; broad lines come from inner regions with exceptionally high density or atypical dust properties; and [Fe II] emission arises from dense gas between broad and narrow-line regions. One of our objects, J1025+1402 (nicknamed The Egg), shows extremely high equivalent width Na D, K I, Fe II, and Ca II triplet absorption lines, suggesting the presence of a cool (5000 K), metal-enriched gas envelope. The optical/near-IR continua of these LRDs are also consistent with theoretical models featuring an atmosphere around black holes. The WISE-detected IR emission is consistent with weak dust emission of K. We propose a conceptual model consisting of a largely thermalized cool-gas envelope surrounding the central black hole and an extended emission line region with high-density outflowing gas to explain the observed properties of these local LRDs.
show]xiaojinglin.astro@gmail.com
show]xfan@arizona.edu
]zcai@tsinghua.edu.cn
1 Introduction
The so-called “little red dots” (LRDs) have been among the most significant discoveries of the early years of JWST operations (e.g., I. Labbe et al., 2025; J. Matthee et al., 2024; X. Lin et al., 2024; H. B. Akins et al., 2024). These sources appear compact even in JWST imaging, and the majority of them (over 70–80%) exhibit broad Balmer emission lines with FWHM km s-1, indicative of black holes (BHs) of – (e.g., J. E. Greene et al., 2024; D. D. Kocevski et al., 2024; J. Zhang et al., 2025). They display V-shaped spectral energy distributions (SEDs), characterized by a blue UV continuum and a red optical continuum (D. D. Kocevski et al., 2024; D. J. Setton et al., 2024; R. E. Hviding et al., 2025). Moreover, a large fraction of these V-shaped objects present Balmer absorption on top of the broad emission line, which is rarely seen in other type-1 AGNs (e.g., J. Matthee et al., 2024; X. Lin et al., 2024; B. Wang et al., 2025a; F. D’Eugenio et al., 2025).
Debates about the nature of LRDs are ongoing, particularly concerning the origin of their V-shaped SEDs, the spectral shape inflection near the Balmer break, and the prevalence of Balmer absorption (e.g., D. J. Setton et al., 2024; X. Lin et al., 2024). Early studies have shown that if the Balmer break arises from massive host galaxies, it would challenge our current understanding of structure formation (e.g., B. Wang et al., 2024; I. Labbe et al., 2024). Alternatively, if these features are attributed to reddened AGNs, it is difficult to account for the lack of strong rest-frame IR detections for most high-redshift LRDs, by JWST/MIRI, Spitzer, or ALMA (e.g., P. G. Pérez-González et al., 2024; C. C. Williams et al., 2024; Y. Ma et al., 2024; D. J. Setton et al., 2025). New models have started to explore the distribution of dust (Z. Li et al., 2025; K. Chen et al., 2025), along with new constraints on the dust content in LRDs from multi-wavelength data (C. M. Casey et al., 2024, 2025).
The unusual continuum shape of LRDs has motivated multiple theoretical models (K. Inayoshi et al., 2025; D. Kido et al., 2025). Recent studies have proposed that BHs embedded in dense gas with a high covering fraction can explain the presence of both the observed Balmer breaks and absorption features (K. Inayoshi & R. Maiolino, 2024). Such models have been applied to several high-redshift LRDs and have achieved reasonably good fits to their spectra (X. Ji et al., 2025; A. J. Taylor et al., 2025a; A. de Graaff et al., 2025; R. P. Naidu et al., 2025). However, for sources with rest-frame IR detections, these models struggle to simultaneously reproduce the weak IR photometry and the shape of Balmer breaks. They typically require either an extremely high column density of hydrogen gas (e.g., ), exotic modified dust attenuation laws, or intrinsically red AGN SEDs (A. de Graaff et al., 2025). Even with these additional free parameters, challenges still remain for fully explaining the IR broadband photometry (e.g., A. de Graaff et al., 2025; R. P. Naidu et al., 2025; A. J. Taylor et al., 2025a).
Currently, only a limited number of high-redshift LRDs have been detected with JWST/MIRI, while the vast majority remain either not covered by or undetected in the mid-IR (e.g., G. C. K. Leung et al., 2024; D. J. Setton et al., 2025; X. Lin et al., 2025; A. J. Taylor et al., 2025a). Most JWST/NIRSpec spectroscopy of LRDs at has been conducted in the prism mode, and high-resolution grating observations of several objects reveal only their brightest features. In contrast, LRDs at low redshifts can be observed with a variety of facilities/instruments at multiple wavelengths, with high spectral and spatial resolution at relatively low cost, providing deep insights into the nature of LRDs as a population. Rest-frame IR and high-resolution grating spectra have been obtained for a small number of luminous LRDs at cosmic noon (e.g., I. Juodžbalis et al., 2024; B. Wang et al., 2025a; A. de Graaff et al., 2025). Systematic searches for LRDs at using wide-field sky surveys are now underway and have already demonstrated successful spectroscopic follow-up with ground-based facilities (R. Lin et al., 2025; Y. Ma et al., 2025; Euclid Collaboration et al., 2025). However, photometrically selected LRDs at cosmic noon are still relatively faint ( mag), making follow-up observations challenging with current ground-based facilities.
Identifying LRDs in the local Universe opens a valuable window into understanding the nature of this population. Recent observations and theoretical predictions suggest that their number density declines significantly at , dropping by 1–2 dex from to (D. D. Kocevski et al., 2024; Y. Ma et al., 2025; K. Inayoshi, 2025). This implies that local LRDs are rare and are linked to physical conditions that may be more common in the early Universe. Their discovery at low redshift would offer critical insights into the evolutionary pathways of LRDs and their connection to BH growth across cosmic time. Moreover, local LRDs may enable spatially resolved observations of themselves, their host galaxies, and their environments, revealing their structure, kinematics, and dust/gas content. These observations would not only be cost-effective but also may unravel details that remain challenging to obtain from their high-redshift counterparts, even with JWST.
In this paper, we present the search and discovery of three local LRDs at =0.1–0.2 from the SDSS spectroscopic database, followed by high signal-to-noise radio ground-based spectroscopic observations. The paper is organized as follows. In §2 we introduce the selection criteria we used to search the SDSS spectral database for local LRDs. We present the SDSS data and follow-up observations on the three selected local LRDs in §3. In §4 we describe their overall properties and spectral features, which are fully consistent with those of JWST-discovered high-redshift LRDs. In §5, we analyze their emission line properties and infer the physical conditions. We then conduct a detailed case study of one object in §6. Finally, in §7 we interpret the observed features to understand their physical nature. Throughout this work, we adopt the AB magnitude system for all photometric measurements. All equivalent widths (EWs) are reported in the rest frame. A flat CDM cosmology is assumed, with , , and .
2 Selection and sample from SDSS
In the literature, a wide range of selection criteria have been used to define the LRD population (e.g., K. N. Hainline et al., 2025; R. E. Hviding et al., 2025). In this paper, we adopt a narrow definition of LRDs, referring specifically to objects that simultaneously satisfy the following criteria: (1) presence of broad Balmer emission lines; (2) compact morphology in the rest-frame optical, particularly in photometric bands dominated by H emission; and (3) a V-shaped SED, characterized by a blue UV continuum slope and red optical continuum slope. A detailed description of the selection of local LRDs will be presented in an upcoming paper. Here, we provide a summary of the procedure.
2.1 Selection
We first built a library of high-redshift LRDs with well-characterized UV-to-optical spectra. We adopted the compilation of high-redshift LRDs from D. J. Setton et al. (2024), which exhibit clear V-shaped SEDs with inflection points near the Balmer limit. Their JWST/NIRSpec PRISM spectra were obtained from the DAWN JWST Archive 111https://dawn-cph.github.io/dja/.
We then searched for local LRD candidates in the SDSS DR17222https://www.sdss4.org/dr17/ spectroscopic database (K. N. Abazajian et al., 2009; Abdurro’uf et al., 2022). We started with the GUVmatch333https://archive.stsci.edu/hlsp/guvcat catalog, which cross-matches GALEX far-UV (FUV) and near-UV photometry (NUV) with SDSS DR14 ugriz photometry. We shifted the high-redshift LRD PRISM spectra to –0.5 to serve as templates for our local LRD search. This redshift range was chosen to ensure that the H emission line falls within the SDSS spectroscopic wavelength coverage. We then re-projected the template spectra to GALEX and SDSS photometry to obtain the expected broadband photometry of local LRDs. Using these synthetic colors, we searched for candidates in GUVmatch with similar SED shapes. Specifically, we computed the Euclidean distance in magnitude space between the observed photometry and each individual template. This distance was defined as the root mean square of the magnitude differences across the GALEX and SDSS bands, normalized by the -band flux. Then, we selected sources with a distance of less than one magnitude from any of the templates. Finally, we imposed a compactness criterion by requiring the SDSS-reported Petrosian radii (V. Petrosian, 1976) in both the g and r bands to be smaller than 1.8 arcsec (physical scale of 3.3 kpc at ). This morphological criterion was relatively permissive, and variations in the threshold do not affect the selection results.
After the photometric selection, we retrieved the SDSS spectra of the selected candidates. We examined their emission line properties using the SPZLINE extension, which provides automated emission-line fits from the SDSS pipeline (A. S. Bolton et al., 2012). We first required the EWs of [O III] 5008 to exceed 10 Å, and the line ratio [O III] 5008/[O II] 3727 (O32) to be greater than 10. The first criterion effectively excludes stellar objects, quiescent galaxies, and post-starburst galaxies, which could exhibit strong Balmer breaks and are prevalent in the local Universe. The second criterion ensures that the selected galaxies have physical conditions similar to those of high-redshift galaxies.
To further remove contaminants and confirm the presence of red optical continua, we inspected the continuum underlying emission lines in the SPZLINE extension. We required the continuum under H to be redder than that under [O III] 5008, i.e., in space. We also required the [O III] 5008 continuum to be redder than that under [O II] 3727.
Although we define LRDs by the presence of broad H lines, up to this stage, the broad-line criterion has not yet been explicitly applied during the selection process above. Nonetheless, all targets selected based on the above criteria have already exhibited clear broad H emission in their SDSS spectra. We therefore obtained a sample of local sources that meet all the criteria defining LRDs from the SDSS DR17 spectroscopic database.
2.2 Sample from SDSS
In this paper, we report the discoveries and follow-up observations of three local LRDs: SDSS J102530.29+140207.3 (J1025+1402 hereafter), SDSS J104755.92+073951.2 (J1047+0739), and SDSS J102208.52+084156.1 (J1022+0841). The full sample will be presented in a forthcoming paper.
Figure 1 shows the images, photometry, and spectra of the three selected local LRDs. These objects closely match their corresponding LRD templates and exhibit unambiguous V-shaped UV-to-optical SEDs.
In fact, the local LRDs presented in this paper were originally selected as quasar candidates (G. T. Richards et al., 2002). The SDSS target selection flag for J1025+1402 and J1022+0841 is QSO_HIZ, as their colors meet the criteria for selecting high-redshift quasars, primarily at . Both objects show rising flux toward and peaking in the band, followed by a decline toward the bands, resembling the spectral shape of a Lyman break with a blue UV continuum, characteristic of quasars at . J1047+0739 is flagged as both QSO_CAP and QSO_HIZ: the former because its color meets the selection criteria for quasars at , and the latter because it satisfies a relaxed criterion for high-redshift quasars. These flags indicate that the three selected objects are outliers in both the and color spaces, making them easily misclassified during color selection.
Our LRD selection is incomplete, since it requires that the object has an existing SDSS spectrum. The SDSS main galaxy survey is biased against LRDs because of their compact morphology, which is one of the reasons that they are labeled as quasar candidates. Therefore, we estimate a conservative lower limit on the number density of local LRDs at =0–0.5, based on the effective SDSS survey area of 9,376 deg2. The discovery of these three objects yields an estimated lower limit of approximately within this redshift range. Our estimated number density at =0–0.5 is more than several orders of magnitude higher than the extrapolation of the log-normal distribution fitted to the LRD population at 4–7 (K. Inayoshi, 2025). However, this large discrepancy arises in the exponentially damped tail of the assumed distribution, where such divergence is expected due to the steep functional form and limited constraints at low redshifts. We find that a modest increase in the dispersion (by 10–20%) of the log-normal model from K. Inayoshi (2025) yields a redshift evolution of the LRD abundance consistent not only with measurements at 2–3 (Y. Ma et al., 2025), but also with our observations at =0–0.5. Nonetheless, this trend suggests that the BH accretion mode giving rise to LRDs, while rare in the local Universe, is not exclusively unique to the early Universe.
From the SDSS spectra, J1025+1402 shows clear blueshifted H absorption, while J1022+0841 exhibits tentative but unresolved H absorption. Both features are confirmed by our spectroscopic follow-up with higher spectral resolution (see §4.3). J1025+1402 and J1047+0739 were identified by Y. I. Izotov & T. X. Thuan (2008) as metal-poor AGNs in SDSS DR5, with 12 + (O/H) = 7.36 and 7.99, respectively. C. Simmonds et al. (2016) further noted that they are weak in X-rays, revealed by the non-detection of Chandra. C. J. Burke et al. (2021) found no significant variability in either optical broadband photometry or broad H luminosity.
3 Data and Observations
3.1 Photometry
The three targets were observed by multiple wide-field sky surveys. We adopt photometric data from the following surveys: FUV and NUV from GALEX DR6 (L. Bianchi & B. Shiao, 2020), ugriz Petrosian photometry from SDSS DR17 (K. N. Abazajian et al., 2009; Abdurro’uf et al., 2022), PSF photometry from Pan-STARRS DR2 (H. A. Flewelling et al., 2020), mean photometry from GAIA DR3 (Gaia Collaboration et al., 2023), photometry from Legacy Surveys DR10 (A. Dey et al., 2019), Petrosian photometry from UKIDSS DR11PLUS (A. Lawrence et al., 2007), and W1, W2, W3, W4 photometry from WISE (E. L. Wright et al., 2010). The GALEX photometry is taken from the GUVmatch catalog. The WISE photometry has been incorporated into the Legacy Survey catalog based on the unWISE images (E. F. Schlafly et al., 2019). All other data are retrieved from Astro Data Lab444https://datalab.noirlab.edu/ (M. J. Fitzpatrick et al., 2014; R. Nikutta et al., 2020).
3.2 Spectroscopic Follow-up Observations
We obtained high signal-to-noise-ratio (S/N) LBT/MODS and Magellan/FIRE spectroscopy to characterize the UV-to-NIR spectra of the three local LRDs. We also obtained MMT/Binospec spectra of J1025+1402 and J1022+0841 to measure their high-resolution H profiles.
The LBT/MODS (R. W. Pogge et al., 2010) is a pair of dual-channel spectrographs on the 28.4m Large Binocular Telescope, with the blue channel covering 3300–6000 Å and the red channel 5000–10000 Å. We observed the three targets on February 24, 2025, using 1.0-arcsec slits in grating mode under an average seeing of 0.8 arcsec. This setup achieved a spectral resolution of –1500 in the blue channel and –1700 in the red channel. Observations of each target included 3600 s exposures in both the blue and red channels, with the two LBT telescopes observing simultaneously. As a result, each target effectively has 6600 s exposures in both the blue and red channels.
The Magellan/FIRE (R. A. Simcoe et al., 2013) is a near-IR echelle spectrograph covering 0.8–2.5 on the 6.5m Magellan Baade Telescope. We observed J1025+1402 using a 0.75-arcsec slit on December 21, 2024, reaching –6000. The observation included 5900 s exposures. We observed J1047+0739 and J1022+0841 using 1-arcsec slits on March 11, 2025, with 7900 s exposures for J1047+0739 and 6900 s exposures for J1022+0841, reaching –4000. The seeing for these observations was –0.7 arcsec.
The MMT/Binospec observation of J1025+1402 was conducted in the IFU mode (D. Fabricant et al., 2019, 2025) on December 7, 2024, using the 600 line/mm grating. The observation consisted of s exposures and achieved a spectral resolution of . The seeing during the observation was approximately 1.6 arcsec. For J1022+0841, the observation was conducted in long-slit mode on January 23, 2024, using a 1-arcsec slit and the 600 line/mm grating, under a seeing of approximately 2.3 arcsec. The observation included s exposures, with a spectral resolution of .
The LBT/MODS and Magellan/FIRE data were reduced with PypeIt (J. Prochaska et al., 2020). We performed bias subtraction, flat-fielding, wavelength calibration, sky subtraction, spectral extraction, and telluric correction. To correct for slit losses and calibrate the absolute flux of the LBT/MODS spectra, we first matched them to the SDSS spectra using a wavelength-dependent linear function. The spectra in SDSS DR17 have been processed with an updated flux calibration algorithm, yielding significant improvements at wavelengths below 6000 Å555https://www.sdss4.org/dr17/algorithms/spectrophotometry/. We confirmed that the SDSS spectra closely follow the shape of the corresponding photometry. Then, we applied a constant scaling factor to align the semi-calibrated spectra with the SDSS photometry. The Magellan/FIRE spectra were scaled to match the UKIDSS -band photometry using a constant factor. The MMT/Binospec data were reduced using the Binospec pipeline (J. Kansky et al., 2019). The IFU data were extracted with a rectangular aperture matching the seeing disk. Throughout this work, most measurements are based on MODS and FIRE spectra, while Binospec spectra are used exclusively to analyze the H absorber profiles of J1025+1402 and J1022+0841.
The SDSS spectra have a resolution of –2500. We convolve the SDSS spectra to match the spectral resolution of the MODS data. These SDSS spectra reach S/N2.5 per Å at 4000–5000 Å and S/N9 at 7000–8000 Å for J1025+1402. In stark contrast, the MODS-B and MODS-R spectra achieve S/N13 and 81, respectively, improving the S/N by factors of 5 (blue) and 9 (red). For J1047+0739, the S/N improves from 3.5 to 19 at 4000–5000 Å, and from 6 to 60 at 7000–8000 Å. For J1022+0841, the S/N increases from 0.6 to 5.5 at 4000–5000 Å, and from 2.6 to 29 at 7000–8000 Å.
Figure 1 shows the UV-to-NIR spectra, multi-wavelength photometry, and images of the three targets. The LBT/MODS spectra confirm the presence of the V-shape of the UV-to-optical continua. The declining near-IR spectra in space are consistent with the rest-frame IR photometry of bright high-redshift LRDs detected by MIRI (e.g., B. Wang et al., 2025a; A. de Graaff et al., 2025; A. J. Taylor et al., 2025a; D. J. Setton et al., 2025).
4 Evidence as local LRDs
Basic property
Name | RA | DEC | (mag) | 12 + | ||||
( erg s-1) | ( erg s-1) | ( erg s-1) | ||||||
J102530.29+140207.3 | 156.37622 | 14.03586 | 0.1007 | 19.003 | 7.47 0.30 | 4.95 | 7.550.03 | |
J104755.92+073951.2 | 161.98302 | 7.66423 | 0.1682 | 18.846 | 18.5 0.53 | 8.000.02 | ||
J102208.52+084156.1 | 155.53552 | 8.69892 | 0.2227 | 20.011 | 17.4 0.31 | – | 7.430.03 |
H emission-line and BH properties
Name | FWHM | FWHMHα,narrow | |||
( erg s-1) | ( erg s-1) | (km s-1) | (km s-1) | ||
J1025+1402 | 8.68 0.13 | 1.33 0.02 | 983 10 | unresolved | |
J1047+0739 | 27.51 0.58 | 18.86 0.15 | 949 25 | 47 3 | 6.50 0.09 |
J1022+0841 | 48.10 0.49 | 3.43 1.10 | 1022 18 | 6.70 0.09 | |
In this section, we present the observed characteristics of three local objects that closely resemble high-redshift LRDs. The section is organized as follows: morphology and continuum luminosities are discussed in §4.1; broad emission lines and BH masses in §4.2; Balmer and He I absorption features in §4.3; variability properties in §4.4; X-ray and radio properties in §4.5; and a summary of the observational evidence as LRDs is provided in §4.6. We present their zoomed-in spectra in Figures 2, 3, and 4, and summarize the derived properties in Table 1.
4.1 Morphology and Continuum Luminosity
All three objects exhibit very compact morphologies across the optical bands. They are not spatially resolved in LBT/MODS direct imaging with a seeing of 0.9″. This implies that the diameter of J1025+1402 is comparable to or smaller than 1.7 pkpc, while those of J1047+0739 and J1022+0841 are constrained to be smaller than 2.5 pkpc and 2.3 pkpc, respectively.
J1047+0739 was observed with the HST/ACS FR782N ramp filter narrow-band imaging that covers both the H line and the adjacent continuum (GO-15617; PI: F. Bauer), achieving a spatial resolution of 206 pc. The continuum map reveals a bright central point source contributing about 85% of the light, alongside a compact yet clumpy host galaxy with a half-light radius of 530 pc. The H line map is dominated by the central point source with minimal host contribution. The spatially resolved map shows that, in J1047+0739, AGN emission dominates both the continuum and the H line, in agreement with our definition of LRDs, although with a non-negligible contribution of the stellar light from the host galaxy to the continuum. The detailed morphological analysis is presented in Appendix A. Future high-resolution imaging with HST, JWST, or ground-based AO will further constrain the properties of the physical sizes of the local LRDs and their host galaxy properties.
All three objects exhibit continuum shapes characterized by pronounced V-shapes and declining IR slopes. The inflections occur near the Balmer limit. The continuum luminosity at rest-frame 5100Å () is around erg s-1. J1047+0739 exhibits a potential Balmer jump, presumably due to the presence of a host galaxy. For all three objects, the WISE photometry at rest-frame 3–20 µm reveals the presence of weak dust emission, at a wavelength range beyond what is possible with JWST observations of high-redshift LRDs. We integrate the continua from rest-frame 2000 Å to 20 µm to estimate the total continuum luminosity over this wavelength range, yielding – erg s-1 for all three objects.
4.2 Broad Emission Lines and BH masses
The three objects show clear broad profiles in H, H, Pa to Pa, and several He I lines (e.g., He I 7065, 10830). J1025+1402 and J1022+0841 also exhibit broad infrared O I emission lines, consistent with the line broadening originating from gas motions in the BLRs around BHs. The methodology of the line measurements is detailed in Appendix B. Specifically for the H lines, we model the narrow component using a single narrow Gaussian. The broad components are fitted with two to three broad Gaussians. The [N II] 6548, 6583 doublet is modeled with two narrow Gaussians sharing the same FWHM as the narrow H component and fixed line ratios of 3. For J1025+1402 and J1022+0841, additional Gaussian absorbers are applied atop the broad emission lines. The measurements from the MODS spectra are summarized in Table 1. We also report the narrow H FWHMs of J1025+1402 and J1022+0841 from the Binospec spectra, taking advantage of its higher spectral resolution.
All three objects exhibit broad H emission with luminosities around erg s-1, and their individual broad components exhibit FWHMs ranging from km s-1 to 2000 km s-1 , with the composite FWHMs around 1000 km s-1. We estimate their BH masses using the H luminosity (), the composite broad-line H FWHM, and the empirical relation from J. E. Greene & L. C. Ho (2005). We find BH masses around –. This scaling relation for BH mass estimates is widely used in high-redshift LRD studies (e.g., J. Matthee et al., 2024; X. Lin et al., 2024, 2025; A. J. Taylor et al., 2025b). However, this estimator might not be fully applicable to LRDs, if their physical nature is different from classical AGNs, particularly if their line profiles and continua arise from distinct physical mechanisms (see §7). The measured values in these objects are 3–6 times higher than the predicted by the – relation in J. E. Greene & L. C. Ho (2005), where refers to the total (broad + narrow) H luminosity from AGNs. It remains unclear whether or reliably traces the BLR size in these systems.
4.3 Balmer and He I Absorption
Figure 5 presents the line profiles of H, H, He I, and Pa for the three local LRDs. All three objects exhibit absorption lines in their Balmer series or He I 10830 lines. For H, we model the absorber as Gaussian components superimposed on the intrinsic broad H emission. We then simultaneously fit the absorbers, the narrow H, broad H, and the [N II] 6548, 6583 doublet. For H and He I 10830, absorbed regions are masked, and the intrinsic emission line profile is modeled using a narrow Gaussian combined with multiple broad Gaussians. The detailed fitting methodology and complete absorption measurement information are described in Appendix B.
J1025+1402 exhibits a redshifted H absorption feature superimposed on the broad H emission and continuum, with km s-1 and Å. A similar redshifted H absorber is also seen in the triply imaged LRD in the Abell 2744 field (X. Ji et al., 2025). J1025+1402 and J1047+0739 both have H absorption, superimposed on the broad H emission. The H absorber of J1025+1402 has an EW of Å, and is blueshifted by km s-1 relative to the center of the broad H component. The H absorber in J1022+0841 can be described with two Gaussian components, one blueshifted by km s-1 and one redshifted by km s-1. Their EWs are Å and Å, respectively, yielding a total EW of Å. An alternative solution is one Gaussian absorber with km s-1 and EW of Å (see Appendix B). J1047+0739 does not show clear Balmer absorption.
The lack of H absorbers corresponding to H absorbers in J1025+1402 and J1022+0841 might be attributed to the S/N of the H spectra. However, a puzzling aspect of J1025+1402 is that its redshifted H absorber does not have corresponding H absorption. This suggests that the two lines might have distinct origins.
All three targets show prominent blueshifted He I 10830 absorption. The He I 10830 absorber of J1025+1402 and J1047+0739 has multiple components with distinct kinematics (Figure 6). The line profile of the J1025+1402 He I 10830 absorber, as shown in Figure 5, can be described using three Gaussian components, blueshifted with , and km s-1, respectively. This profile may alternatively result from a mixture of saturated He I absorbers. The line profile of the J1047+0739 He I 10830 absorber can be described using five Gaussian components with , and km s-1, respectively. For J1022+0841, the He I 10830 absorber is well described by a single Gaussian component with km s-1. They highlight the prevalence of He I absorption in local LRDs. As shown in Figure 6, for J1025+1402 and J1022+0841, the velocities of the He I and H absorbers are not well aligned, indicating complex outflow kinematics.
For reference, we also show the Pa line profile in Figure 5. No absorption features are associated with Pa. This suggests that the Balmer absorbers originate from hydrogen atoms with a population elevated to the energy level, but not to the level.
4.4 Variability
We examine the variability of all three objects in the optical () and infrared (WISE W1 and W2) bands. For the optical bands, we retrieve multi-epoch PSF-modeled photometry from the Pan-STARRS DR2 detection catalog. Over 1500 days (1300 days in the rest frame), we find no significant variability, with observed variation of 0.1–0.2 mag consistent with the scatter in non-variable sources (T. Simm et al., 2015). In the WISE W1 and W2 bands, we find variations of 0.25–0.5 mag over 5000 days (4000 days in the rest frame), also consistent with the typical scatter expected for non-variable sources. We thus conclude that no significant variability is detected in either the rest-frame optical or infrared bands, consistent with the lack of significant variability seen in high-redshift LRDs (M. Kokubo & Y. Harikane, 2024; Z. Zhang et al., 2025; X. Lin et al., 2025; W. L. Tee et al., 2025).
C. J. Burke et al. (2021) reported no significant variability in the broad H luminosity of J1025+1402 and J1047+0739 over a continuous 15-year period. The H luminosity range of J1025+1402 in C. J. Burke et al. (2021) is approximately – erg s-1, and for J1022+0841 it is – erg s-1, both well consistent with our measurements. They also presented V-band light curves from the Catalina Real-Time Transient Survey (CRTS; A. J. Drake et al., 2009) and -band light curves from the Zwicky Transient Facility (ZTF; F. J. Masci et al., 2019) for both objects. Although the limited quality of the light curves precluded definitive conclusions, they found potential low-level variability consistent with the stochastic damped random walk model of AGN variability (B. C. Kelly et al., 2009; C. L. MacLeod et al., 2010).
4.5 X-ray and Radio
C. Simmonds et al. (2016) reported net counts of Chandra for J1025+1402 in a net exposure time of 4937 s, placing an upper limit666Assuming a power law spectrum absorbed by Galactic . of . For J1047+0739, they reported net counts in 4782 s. We further examined the Chandra Source Catalog Release 2.1 (I. N. Evans et al., 2024), which includes more recent observations of J1047+0739, and found net counts in the 0.5–7 keV band over a 53 ks exposure. This corresponds to an X-ray luminosity of , although the detection is only significant at . For J1022+0841, no Chandra observations are currently available.
The comparison of the X-ray emission of J1025+1402 and J1047+0739 with that of typical type-1 AGNs is detailed in C. Simmonds et al. (2016). The upper limits or X-ray detections suggest that J1025+1402 and J1047+0739 are X-ray weak, exhibiting ratios at least 1–2 dex below those of typical type-1 AGNs with similar broad H luminosities (J. E. Greene & L. C. Ho, 2004; F. Panessa et al., 2006; L.-B. Desroches et al., 2009). This X-ray weakness is consistent with that seen in high-redshift LRDs (M. Yue et al., 2024; T. T. Ananna et al., 2024).
We examine the radio data from the VLA Sky Survey (M. Lacy et al., 2020), the TIFR GMRT Sky Survey (H. T. Intema et al., 2017), and the VLA FIRST Survey (R. H. Becker et al., 1995). We find no radio detections at 153 MHz, 1.4 GHz, and 3 GHz for any of the three sources. The corresponding upper limits for J1025+1402, J1047+0739, and J1022+0841 are: 13.8, 12.5, and 15.1 mJy at 153 MHz; 0.7, 0.7, and 0.8 mJy at 1.4 GHz; and 0.4, 0.5, and 0.4 mJy at 3 GHz, respectively. The upper limits of radio-loudness (, M. Sikora et al. 2007) for the three objects are estimated to be 8, 12, and 20, respectively. The current radio data are not deep enough to determine whether these sources are radio-quiet ().
4.6 Summary of Evidences as LRDs
As discussed above, the three objects at – satisfy all the defining criteria for LRDs. We summarize the similarity between the observed properties of the three local sources to high-redshift LRDs as follows:
compact morphology;
V-shaped continuum (blue UV continuum slope, red optical continuum slope) and decreasing IR continuum at µm;
broad Balmer and Paschen emission lines, consistent with – BHs;
weak X-ray emission;
no significant (or low-level) variability;
high occurrence rate of Balmer and He I 10830 absorption.
Their broad H luminosities, erg s-1, fall within the range observed in high-redshift LRDs. If redshifted to , their expected apparent magnitudes in JWST/NIRCam F444W would be 25.7–26.8 mag. These values lie within the magnitude range observed for the diverse LRDs discovered at (e.g., J. E. Greene et al., 2024; D. D. Kocevski et al., 2024; H. B. Akins et al., 2024; D. D. Kocevski et al., 2025).
5 Emission Line Analysis
In this section, we discuss in detail the properties of the emission lines detected in the local LRDs to constrain their origins, gas density, metallicity, and ionizing sources.
5.1 Balmer Lines and Paschen Line Decrements.
We utilize the rich series of Balmer and Paschen lines to measure the Balmer and Paschen decrements. For high-order hydrogen lines with insufficient S/N to detect reliable broad components, we fit them with a single Gaussian representing the narrow component, assuming that the broad components are buried in the noise. All measurements are based on the MODS spectra for consistency. As shown in Figure 7, the narrow Balmer lines from H to H12 in J1025+1402 and J1022+0841 closely match the intrinsic Case B recombination predictions (, ). The narrow Balmer lines in J1047+0739 are consistent with the SMC dust attenuation law (Y. C. Pei, 1992) with modest attenuation of . The exception is H and H8, which are blended with [Ne III] and He I 3889, respectively. The agreement indicates that the narrow hydrogen lines originate from regions with minimal dust attenuation. However, we caution that the decomposition of narrow and broad components is degenerate, and the strength of the potential underlying absorption remains uncertain.
The narrow Paschen lines generally follow the Case B predictions, similar to the Balmer series, although J1022+0841 exhibits significant scatter due to limited S/N. The P line in J1025+1402 deviates from the Case B prediction, but its flux measurement may be contaminated by nearby He I emission lines adjacent to the Paschen transitions. This issue may also affect the measurement in J1022+0841. In contrast, the He I emission is particularly strong in J1047+0739, allowing both the He I and Paschen lines to be reliably fitted.
The broad H/H line ratios are in all three objects. Such large decrements may be attributed to either radiative transfer effects in high-density gas (e.g., collisional excitation and de-excitation, resonance scattering, etc.) or substantial dust attenuation. Notably, the measured decrements, both in the broad and narrow components, are well consistent with those observed in high-redshift LRDs (M. Brooks et al., 2025). These values are comparable to those expected for the SMC attenuation law with under Case B conditions. However, one caveat in attributing the broad-line decrements to dust is the survival of dust grains in the BLRs, which likely lies within the sublimation zone (e.g., P. G. Martin & G. J. Ferland, 1980; M. Kishimoto et al., 2007; C. M. Gaskell & P. Z. Harrington, 2018; A. Baskin & A. Laor, 2018). For comparison, type-1 AGNs in the SDSS typically show broad H/H ratios around 3.7, with the most extreme cases reaching 5 (D. Ilić et al., 2012). The much larger broad Balmer decrements observed in LRDs suggest that their BLRs may have extreme conditions with unusually dense gas or atypical dust properties, which differ substantially from those in typical type-1 AGNs (e.g., K. Inayoshi et al., 2022).
5.2 BPT Diagram and Metallicity
We estimate the BPT diagnostics (J. A. Baldwin et al., 1981) and the gas-phase metallicity of the narrow line regions (NLRs). To measure the fluxes of forbidden lines, we fit Gaussian profiles. Since the [O III] and lines are blended with [Fe II] emission, their fluxes are measured from spectra where the [Fe II] contribution has been subtracted using the best-fit models (see §5.3).
We show the positions of the three local LRDs on the [N II], [S II], and [O I] BPT diagrams in Figure 8. In the [N II] and [S II] diagrams, all three LRDs lie near the boundary between H II regions and AGN ionization. In the [O I] diagram, J1025+1402 and J1022+0841 fall within the Seyfert region, while J1047+0739 lies near the edge of the H II region. Their positions on the [N II] and [S II] diagrams also overlap with those of low-metallicity AGNs (L. J. Kewley et al., 2013; C. Simmonds et al., 2016). Notably, the locations of these local LRDs closely match that of the cosmic-noon LRD in I. Juodžbalis et al. (2024). This suggests that the narrow emission lines in both local and high-redshift LRDs are excited by similar mechanisms associated with AGN activity.
Additionally, we detect [Fe V] 4071 emission lines in all three objects. This high-ionization line further indicates the presence of a hard ionizing radiation field and AGN activity.
We determine the metallicities of the NLRs using the direct electron temperature () method. The of the O++ zone is derived from the ratio between [O III] and [O III], assuming cm-3. The estimated (O++) is K for J1025+1402, K for J1047+0739, and K for J1022+0841. The of the O+ zone is estimated using the empirical (O+)–(O++) relation from Y. I. Izotov et al. (2006). We apply a dust attenuation correction of to J1047+0739. As summarized in Table 1, all three objects exhibit metal-poor NLRs with 12 + ranging from 7.42 to 8.00. The values are broadly consistent with those reported by Y. I. Izotov & T. X. Thuan (2008) for J1025+1402 and J1047+0739.
We note that although the [O II] lines are detected in J1047+0739, they are significantly affected by recombination excitation enhanced by the hard ionizing radiation field (X. W. Liu et al., 2000; S. Tan et al., 2024). Therefore, we do not use them for (O+) diagnostics. The (O+) value directly derived from the [O II] and [O II] line ratios exceeds K, approximately twice that of (O++). This result further suggests an AGN origin for the narrow emission lines driven by hard ionization.
Based on the Balmer decrements, BPT diagram, and direct -based metallicity, we conclude that the narrow lines in local LRDs originate from metal-poor NLRs powered by AGNs with minimal dust content.
5.3 [Fe II] Emission Analysis
The series of narrow [Fe II] emission lines observed in J1025+1402 and J1022+0841 stands in stark contrast to the broad Fe emission lines typically observed in type-1 AGNs (e.g., I Zw 1, M. P. Véron-Cetty et al., 2004). Though narrow optical [Fe II] lines have been reported in type-1 AGNs, they are blended with broad Fe lines and exhibit transitions and line ratios different from those in our sample; these narrow Fe lines are absent in type-2 AGNs (X.-B. Dong et al., 2010). To constrain the physical conditions such as temperature and density, we employ cloudy for photoionization modeling (C23, C. M. Gunasekera et al., 2023), with the latest Fe II atomic database (R. T. Smyth et al., 2019). We consider scenarios in which the gas is photoionized by AGN radiation. The modeling methodology is detailed in Appendix C.
The AGN photoionization models yield best-fit parameters for both J1025+1402 and J1022+0841 of , , and . The relative [Fe II] line ratios are not strongly sensitive to metallicity, so we adopt a fiducial value of without attempting to constrain it. The average temperature of the [Fe II]-emitting gas is approximately K, ranging from K in the inner region to K on the surface. As shown in Figure 9, the best-fit model reproduces the observed [Fe II] line ratios of J1025+1402 reasonably well, although some lines (e.g., [Fe II] 5278) remain imperfectly matched. A similar level of agreement is achieved for J1022+0841 using the same best-fit parameters (Figure C.1).
The inferred density lies between typical values for the BLRs and NLRs. The FWHMs of the [Fe II] lines are also intermediate ( km s-1 for J1025+1402 and km s-1 for J1022+0841), broader than the narrow H and [O III] (H –70 km s-1 in Binospec, [O III] not resolved in MODS). This suggests that the [Fe II] emission may arise from dense gas located outside the BLRs but inside the NLRs. We emphasize that our models are qualitative. Fe emission lines in AGNs are inherently complex and can be influenced by a variety of physical processes, making comprehensive theoretical modeling challenging (e.g., A. Sarkar et al., 2021; A. Pandey et al., 2025). Further refinement is required to fully understand the origin and nature of [Fe II] emission in the LRD population.
6 Case study: J1025+1402, “The Egg”
In this section, we connect the observed spectral characteristics to the underlying physical picture of LRDs, using J1025+1402 as a case study. J1025+1402 is the most extreme object in our sample, distinguished by its prominently V-shaped continuum and high-EW absorption features in the optical spectrum. These characteristics provide important insights into the nature of LRDs. We extend the insights gained from J1025+1402 to the broader LRD population in §7. Although the extreme metal absorbers seen in J1025+1402 are not observed in the other two objects, we attribute this diversity to the effects of geometry, viewing angle and metallicity (see §7).
Hereafter, we nickname J1025+1402 as the Egg, inspired by its Ca-rich envelope, a temperature comparable to that of yellow supergiant stars, and its distinctive geometry. These features will be discussed in more detail below. An independent discovery and analysis of this object is also presented in Ji et al. (in prep).
6.1 Strong Metal Absorption: Evidence for a Cool Gas Envelope
The Egg exhibits exceptionally high-EW Na D 5890, 5896, K I 7665, 7699, Fe II 6148, 6244, 6456, and Ca triplets (8498, 8542, 8662; Ca T) absorption features in its optical continuum. We show their profiles in Figure 10 and summarize their properties in Table 2.
Absorber | EW | FWHM |
(Å) | (km s-1) | |
Na D doublets | ||
K I doublets | ||
Fe II triplets | ||
Ca II triplets |
6.1.1 Na D and K I Absorption
The Na D absorption in The Egg has an EW of Å and K I absorption shows an EW of Å. These two neutral alkali metals can be easily ionized by photons with low ionization potentials (5.14 eV for Na and 4.34 eV for K). This implies that the absorbing gas must be both cool and dense to preserve its neutral state. One possible origin for these absorbers is the cold, self-shielded intervening gas in the interstellar medium (ISM), with a minor contribution from late-type stars. However, the observed Na D EW in The Egg exceeds those typically found in such environments. The Na D EW in The Egg is higher than that observed in most galaxies and AGNs (typically Å; Y.-M. Chen et al., 2010; S. Cazzoli et al., 2016; D. S. N. Rupke et al., 2021; Y. Sun et al., 2024), and is only comparable to a few extreme IR-luminous LINERs (D. S. Rupke et al., 2005). In local galaxies and AGNs, high-EW Na D absorption typically traces gas outflows or inflows with velocities of several hundred km s-1 (S. Cazzoli et al., 2016; G. W. Roberts-Borsani & A. Saintonge, 2019; D. S. N. Rupke et al., 2021; Y. Sun et al., 2024), whereas in The Egg we do not observe significant velocity shifts.
Likewise, the K I EW is orders of magnitude larger than those found in the Galactic interstellar medium or in circumstellar gas clouds around young stellar objects, where values typically range from a few up to 200 mÅ (L. M. Hobbs, 1975; D. E. Welty & L. M. Hobbs, 2001; I. Pascucci et al., 2015). All of these lines of evidence point to a non-stellar and non-ISM origin.
6.1.2 Fe II Absorption
The optical Fe II absorption lines at 6148, 6244, and 6456 are rarely observed in galaxies or AGNs. These features have been reported in a Type Ia supernova (K. Maguire et al., 2023), but with large blueshifts of several thousand km s-1 and broad line widths. In contrast, the Fe II absorption seen in The Egg shows no significant velocity shift and narrow FWHMs of 400 km s-1. Although the origin of these Fe II absorption features remains uncertain, their co-occurrence with Na D, K I, and Ca T absorption may suggest a shared origin.
6.1.3 Ca T Absorption
The Ca T absorption at 8498, 8542 and 8662 Å in The Egg has an EW of Å. In Figure 11, we compare this value to typical Ca T EWs observed in stars, star-forming galaxies, and AGNs. The Ca T EW in The Egg exceeds those of most stars and all known galaxies and AGNs, where Ca T EWs are generally below 10 Å.
Only a few individual stars, such as yellow hypergiants or supergiants (S. V. Mallik, 1997; A. J. Cenarro et al., 2001), can exhibit comparable Ca T EWs. However, reproducing the observed Ca T EWs in the integrated galaxy light would require an unusually exotic stellar population composed solely of these stars.
6.1.4 The Origin: a Cool Gas Envelope around BHs
The unusually high-EW Na D, K I, Fe II, and Ca T absorption lines of The Egg all point to a non-stellar and non-ISM origin. If the optical continuum is AGN-dominated, these absorption lines likely arise from an atmosphere-like gas envelope in the vicinity of the central BHs. Although a stellar population struggles to explain the lines observed in The Egg, the microphysics of line formation may apply to both stellar atmospheres and LRDs. Therefore, a comparison to stellar spectra enables us to infer the physical conditions of the absorbing gas.
In the following, we investigate the EW of Ca T absorption as a key tracer. Figure 12 places The Egg’s CaT absorption in the context of stellar atmospheres. As shown in the left panel of Figure 12, within the stellar catalog presented here, only two metal-rich stars exhibit Ca T EWs greater than that of The Egg. These two stars are classified as an Ia yellow supergiant and a 0–Ia yellow hypergiant, with effective temperature () of 5500 and 5727 K, surface gravities () of 0.0 and 0.4, and metallicities [Fe/H] of and , respectively.
The yellow supergiant and hypergiant stars exhibit the strongest Ca T absorption due to their low surface gravity and their effective temperatures of K. This temperature allows Ca in the atmosphere to exist mainly in the singly-ionized state. In stellar atmospheres, the EWs of Ca T absorption strongly anti-correlate with surface gravity (e.g., S. V. Mallik, 1997; A. J. Cenarro et al., 2002). This is due to the gas opacity changing with the photosphere density, which generally increases with in stellar atmosphere models. As decreases, the photosphere density decreases, leading to a diminished continuum absorption opacity (the main source being H- ions at and at wavelengths around – Å), while the line opacity of Ca II changes only mildly (U. G. Jorgensen et al., 1992). This enhances the contrast between the line and continuum opacities, increasing the EWs of the Ca T absorption. Thus, the Ca T EWs are essentially atmospheric density indicators at a given effective temperature.
We interpret the Ca T absorption in The Egg as originating from an atmosphere-like gas envelope surrounding the BH, with a structure and opacity analogous to that of cool stellar atmospheres. We apply stellar Ca T diagnostics to estimate to first order. The right panel of Figure 12 demonstrates the stellar Ca T EWs as a function of . The Ca T EW– relation also depends on stellar metallicity. Extrapolating the empirical Ca T EW- relation for metal-rich stars with [Fe/H]0.25 to match the observed Ca T EW of The Egg yields an estimated of approximately . If we assume a slightly lower metallicity ( to ), the inferred would be about . The value may be even lower given the low metallicity we have inferred from the NLR (Section 5.2).
In conclusion, the Ca T absorption in The Egg indicates the presence of a metal-rich cool gas envelope with low surface gravity (). Comparison with known supergiant and hypergiant stars suggests that the of such an atmosphere is likely around 5000–6000 K. However, we emphasize that this is a first-order approximation with simplified assumptions; the actual physical conditions of the absorbing gas may differ substantially from those in stellar atmospheres. In particular, the mapping between and the stellar atmosphere density (and hence opacity) is based on hydrostatic equilibrium, which is not necessarily true for the envelope of The Egg. Nevertheless, as we argue above, the line formation process mainly depends on the atmospheric temperature and density. The inferred value of above may not reflect the dynamical condition of the envelope, but does suggest that its atmosphere density may resemble that of very low-surface gravity stars. Standard stellar photosphere relations imply that at K corresponds to a photosphere density of (e.g., C. J. Hansen et al., 2004, their Equations (4.56)(4.65)). Thus, we expect the photosphere density of The Egg to be on the order of or lower than .
We note that the FWHMs of these metal absorbers are about 300–450 km s-1 (Table 2), significantly broader than those typically observed in stellar atmospheres. For instance, Ca T absorption lines in stars are intrinsically narrow, with FWHMs reaching only up to 70 km s-1 in supergiant stars (A. V. Filippenko & L. C. Ho, 2003). The broad Ca T absorption in The Egg suggests that the cool gas envelope is turbulent or dynamically complex. A more detailed analysis, incorporating tailored modeling of the gas properties, is necessary for a comprehensive understanding.
We have limited our analysis to the Ca T lines; similar atmosphere conditions are presumably responsible for the other absorption lines identified here. However, we do not perform the same analysis on them due to our unawareness of dedicated stellar datasets on these lines. Future characterization of observed stellar spectral lines combined with analysis of theoretical spectral libraries can directly test our scenario.
6.2 Optical to IR Continuum: Evidence of Thermalized Emission
As discussed in §6.1, we find evidence for cool gas envelopes that may be analogous to (supergiant or hypergiant) stellar atmospheres at 5000 K. In this section, we investigate the potential application of atmosphere-like gas envelopes to explain the observed continuum shape of the local LRDs. A similar comparison conducted by A. de Graaff et al. (2025) also noted a resemblance to supergiant stars, though with notable discrepancies. In Figure 13, we compare the spectrum of The Egg to a 5000 K blackbody. The Å regime of The Egg’s spectrum can be reasonably well described by the Rayleigh-Jeans tail of a 5000 K blackbody. However, the Å region shows a significant flux deficit relative to the blue side of the blackbody spectrum, even without subtracting the extrapolation from the UV continuum as a separate component. Varying the blackbody temperature does not yield a better match, as the optical-to-near-IR continuum of The Egg exhibits stronger curvature than a standard blackbody. Such a pseudo-blackbody continuum shape signifies thermalized emission, but may involve intricate physical mechanisms.
Recent theoretical models propose that envelope-like structures can form around SMBHs under super-Eddington accretion and emit thermal radiation with characteristic temperatures of K (D. Kido et al., 2025; H. Liu et al., 2025). To investigate such scenarios, we compare the observed continuum of The Egg with the models from H. Liu et al. (2025), which assume an atmosphere formed by quasi-spherical super-Eddington accretion flows. As shown in Figure 13, one such model from H. Liu et al. (2025) provides a good match to the optical continuum shape of The Egg. This model has a total luminosity and a density of at the photosphere 777 in their notation.; a star at would have this photosphere density at (C. J. Hansen et al., 2004). The photosphere (defined by the Rosseland-mean optical depth) is located at 510 AU with K. The frequency-dependent gas optical depth effects suppress the UV-to-optical flux, making the predicted continuum deviate from a blackbody 888We invoke no dust reddening to the model curve; we have checked that adding a modest level of dust extinction ( mag using the SMC dust extinction law) would further improve the agreement.. Essentially, the optical spectrum is “reddened” by the atmospheric gas; otherwise, we would have to invoke a very steep dust extinction law to explain the optical redness while retaining the declining IR slope (see Appendix D). However, we caution that the model assumes super-Eddington accretion, while current observations do not provide direct evidence of it. The integrated corresponds to about 22% of the Eddington luminosity, based on the listed in Table 1. The true Eddington ratio is unconstrained, as both and the intrinsic bolometric luminosity are uncertain. In addition to the theoretical model, we present an alternative empirical parameterization of the pseudo-blackbody continuum shape in Appendix D.
The origin of the blue UV continuum remains uncertain. Several observations suggest a stellar origin from young host galaxies, based on the lack of UV variability, the absence of high-ionization UV lines, and the extended, complex UV morphologies in a number of high-redshift LRDs (P. Rinaldi et al., 2024; H. B. Akins et al., 2025; W. L. Tee et al., 2025; R. P. Naidu et al., 2025). However, these features do not definitively rule out an AGN origin, either directly from the accretion disk or scattered by electrons/dust (E. Lambrides et al., 2024; Z. Li et al., 2025; K. Inayoshi et al., 2025). This is especially the case when considering the recent detection of AGN-powered high-ionization UV lines in a subset of LRDs (M. Tang et al., 2025). Recently, P. Rinaldi et al. (in prep.) found that a V-shaped continuum similar to that of LRDs originates from the nuclei of a cosmic-noon IR-bright galaxy, further suggesting that AGN emission could contribute to the blue UV continua. At present, we lack spectral information at rest-frame wavelengths below 2000 Å to map diagnostic UV lines that could distinguish between AGN and star-forming activity. As such, the origin of this component remains inconclusive given the current data. Deep UV spectroscopy or polarization measurements, both of which could trace AGN light, is needed to clarify its nature.
6.3 The Dust Emission
The WISE photometry of local LRDs suggests that the dust content in these objects, while not as strong as implied by a standard AGN dust torus, is also not negligible. A typical AGN dust torus, when scaled to LRD luminosities, would show rapidly increasing emission beyond 1–2 µm, peaking around 10 µm, which is clearly not the case for LRDs (M. Stalevski et al., 2016; B. Wang et al., 2025b). All three local LRDs show flattening or an increase in flux beyond rest-frame 10 µm (W3 and W4 bands; Figure 1). At , this wavelength range is beyond JWST coverage, and it is either only partially covered by Spitzer and Herschel or falls into gaps between their coverage, leaving it largely unconstrained for high-redshift LRDs (H. B. Akins et al., 2024; C. M. Casey et al., 2024, 2025; D. J. Setton et al., 2025).
Qualitatively, the IR excess from the WISE photometry in these local LRDs (Figure 1) reveals at least dust components with temperatures approximately between and K. The emission at rest-frame 3 µm corresponds to the peak of a K blackbody, and the emission at 10 µm matches a K blackbody. As shown in Figure 13, if assuming templates from J. Lyu & G. H. Rieke (2021) that include dust components at , 1038, 300, and 90 K (see Appendix §D.2), the IR emission of The Egg can be well reproduced by significant dust components at 1038 K and 300 K. However, the derived dust properties depend sensitively on the chosen templates. We discuss alternative dust templates based on Haro 11 (J. Lyu et al., 2016) and the degeneracies in Appendix D.3. In the absence of longer-wavelength data, it is difficult to robustly constrain the dust properties, particularly the cold dust content with temperatures below 100 K.
7 A conceptual picture of LRDs
Physical Process | Observational Evidence | Caveat and prospects |
BH accretion | Broad emission lines (e.g., Balmer and Paschen series, infrared O I, infrared Fe II) with FWHM 1000 km s-1 | Optically thick gas envelope prevents direct observations of BH accretion process. |
AGN Emission Line Regions | Narrow and broad line regions consistent with AGN ionization; [Fe II] emission lines with FWHM and inferred density between broad and narrow components | Modeling of [Fe II] is complex and potentially degenerate |
Cool gas envelope | Metal absorption lines, especially with extremely high-EW Na D, K I, Fe II, and Ca T absorption features in The Egg; the thermalized emission in good agreement with theoretical atmosphere models (H. Liu et al., 2025) | Requires detailed modeling of opacity, cooling, and circum-BH environments; must also account for the FWHMs of absorption lines |
Non-spherical geometry and viewing angle | Optically thin channels within the envelope are required for AGN ionizing photons to escape and illuminate the BLRs, intermediate [Fe II] regions, and NLRs; may also explain the diversity in continuum shapes and the presence or absence of high-EW metal absorbers. | The impact of viewing angle is difficult to quantify without a statistically large sample. |
Outflowing optically-thick gas | Blueshifted Balmer and He I absorption lines | Their correlation with other properties of LRDs (i.e., Balmer break strength) remains uncertain; the origin of He I emission (ionized gas versus dense neutral gas) requires further study. |
Dust emission | near-IR excess in WISE photometry (rest-frame 3-20 µm) | Constraints on dust components depend on model assumptions; cannot fully characterize dust properties from the AGN to the ISM due to the lack of longer-wavelength data. |
Based on the observational evidence and physical interpretations discussed above, we construct a conceptual picture of LRDs, as shown in Figure 14. Table 3 summarizes the proposed physical processes, their observational signatures, and the associated caveats in our interpretation.
We first conduct an order-of-magnitude estimate of the characteristic physical scales that support the proposed framework. Assuming that the observed emission line broadening is gravitational in origin, the enclosed mass is given by . For a broad emission line with FWHM of 1000 km s-1, and adopting a low surface gravity atmosphere with as implied by the strong Ca T absorption, we derive
In the calculation above, we use the broad emission lines as velocity indicators, rather than the absorption line widths, and place the BLRs near the edge of the photosphere. The kinematics of the BLR clouds are assumed to be primarily governed by gravity, allowing for a reasonably simplified estimate based on Kepler’s laws. In contrast, the broadening of metal absorption lines in the atmosphere is more complex, influenced by intricate processes such as anisotropically distributed turbulence, making their kinematics more difficult to interpret.
The total luminosity of a blackbody radiation with a temperature of 5000 K (Figure 13) and a radius of 700 AU is about erg s-1. This estimated luminosity is consistent within an order of magnitude with the observed luminosities (, Table 1) and the luminosities of the theoretical models (Figure 13 and D.1) for the three local LRDs. For comparison, the corresponding Eddington luminosity for a 8 BH is 1044 erg s-1. The BH mass estimated here is up to an order of magnitude smaller than the H-based BH masses listed in Table 1. Given the approximate nature of these estimates, the overall agreement between the observed properties and those inferred from this conceptual model supports the feasibility of an atmosphere-enshrouded BH model. The individual components of our conceptual framework are described in detail below.
BLRs and NLRs. The broad Balmer and Paschen lines, along with the infrared O I emission lines (in The Egg, see Figure 2), indicate that the observed line broadening arises from gas motions within the BLRs surrounding the BHs. Analysis of the Balmer decrements, BPT diagrams, and metallicity suggests that the narrow emission lines originate from AGN-powered NLRs with low metallicity and minimal dust.
[Fe II] emitting gas intermediate between the BLRs and NLRs. The AGN-excited [Fe II] emission lines indicate gas with a density of approximately cm-3, intermediate between the BLRs and NLRs. The FWHM of the [Fe II] lines ( 200 km s-1 for The Egg and 300 km s-1 for J1022+0841 in MODS spectra) is likewise between that of the broad ( km s-1) and narrow lines (17 km s-1 for The Egg and 77 km s-1 for J1022+0841 in the Binospec spectra). The absence of these [Fe II] lines in J1047+0739 may be attributed to a lower covering fraction of the intermediate emitting regions.
A cool gas envelope with 5000–6000 K and photosphere density of (). The optical–IR continua of the three local LRDs exhibit pseudo-blackbody shapes indicative of thermalized emission and are in good agreement with theoretical atmosphere models (H. Liu et al., 2025). The presence of extremely high-EW metal absorbers (Na D, K I, Fe II, Ca T) in The Egg further indicates a cool, optically thick gas envelope surrounding the BH. Analysis of the Ca T EW, by analogy with stellar atmospheres, suggests an atmospheric temperature of approximately 5000–6000 K and surface gravity of at the photosphere, corresponding to a density of . The size of this envelope is on the order of AU as inferred from the optical-IR luminosity and the effective temperature. A consequence of an optically thick cool gas envelope is that the BH accretion process, i.e., the emission from the accretion disk itself, is not directly observable.
Non-spherical geometry and role of the viewing angle. The observed AGN-powered emission lines imply that the BLRs, intermediate regions, and NLRs cannot be located entirely outside the gas envelope. Otherwise, the –6000 K central source would not produce enough ionizing photons responsible for the recombination lines or the photo-ionization conditions required for forbidden lines (e.g., [Fe II], [N II], O I). Therefore, there should be a relatively optically thin channel that allows high-ionization photons to escape from the disk and gas envelope. We attribute such an optically thin channel along the direction of the line-emitting regions. This non-spherical geometry allows emission lines to form while preserving the envelope structure.
The viewing angle in this non-spherical geometry can explain the diversity in continuum shapes and the presence or absence of metal absorption features among different objects. Changes in viewing angle can modulate the relative contributions of thermalized emission and AGN photoionized radiation. It also determines whether the line of sight intersects the metal-absorbing layers within the cool gas envelope. For example, the line of sight toward The Egg may intersect the optically thick part of the envelope with high metal-line opacity. This orientation results in a significant contribution to the continuum from the thermalized emission, making the metal absorbers observable. J1022+0841 could be oriented closer to the optically thin region, where the optical depth of metal lines is lower. J1047+0739 could lie even closer to the optically thin region, leading to a reduced contribution from the thermalized emission to the total optical continuum. However, we emphasize that this is a speculative scenario, and a larger statistical sample is needed to better understand the roles of geometry and viewing angle. Differences in the metallicities of the cool gas envelop would also contribute to the diversity of observed metal absorption features.
Optically thick outflows. In most high-redshift LRDs, the Balmer absorption lines are blueshifted, with only a few cases showing redshifted H absorption with minimal velocity offsets (X. Lin et al., 2024; X. Ji et al., 2025; F. D’Eugenio et al., 2025). Up to now, all detected He I absorbers in LRDs are blueshifted (I. Juodžbalis et al., 2024; B. Wang et al., 2025a; F. Loiacono et al., 2025). The Balmer absorbers are superimposed on broad Balmer emission lines, and the He I absorbers are superimposed on both the broad emission and the underlying continuum. We thus attribute these absorbers to optically thick outflowing gas located either within or outside the BLRs. The rare instances of redshifted Balmer absorption with small velocity shifts may indicate gas cooling and inflowing back toward the nucleus. Although the relationship between the absorption and LRD properties (e.g., Balmer break strength) remains unclear, it is a plausible connection that merits further exploration.
Dust at – K. The WISE photometry of the three local LRDs is consistent with dust components with temperatures ranging from to K. These dust components, cooler than typical AGN-heated hot dust but warmer than ISM dust, is likely located outside the 5000 K atmosphere yet still within the AGN vicinity. However, we note that the analysis of the dust components is highly uncertain, model-dependent, and subject to degeneracies.
In the proposed scenario, the BH resides in an environment distinct from that of typical type-1 AGNs. As discussed in §4, it remains unclear whether the empirical BH mass estimators calibrated for type-1 AGNs are valid for LRDs, given their potentially different physical conditions. Assuming a BH mass of –, the corresponding Eddington luminosity is – erg s-1. For the three objects, if we take the total integrated luminosity , assuming that the UV continuum is AGN light, the Eddington ratio would be . The current observations do not provide sufficient evidence to determine whether LRDs are in a super-Eddington accretion mode, as proposed by some theoretical scenarios (e.g., K. Inayoshi et al., 2024; F. Pacucci & R. Narayan, 2024). The accretion rate of the central BH and its disk can significantly influence the geometry of both the disk and the surrounding gas envelope (e.g., C. Dotan & N. J. Shaviv, 2011). We leave a detailed exploration of this to future work.
8 Summary
We searched the SDSS spectroscopic database for local LRDs using high-redshift LRD templates and identified three LRDs at –. We followed up these sources with MMT/Binospec, LBT/MODS, and Magellan/FIRE spectroscopy, which reveal detailed spectral features that probe the physical nature of these local LRDs. Our findings are summarized below.
The three local objects closely resemble the high-redshift LRDs. They display compact morphologies. HST narrow-band imaging of J1047+0739 reveals that both the continuum and the H line emission are AGN-dominated, with the compact yet clumpy host galaxy (half-light radius about 530 pc) contributing a minor fraction to the continuum. All the three objects exhibit clear V-shaped UV-to-optical continua with inflections around the Balmer break, and a decreasing IR continuum from Å to 2 µm. All of them show broad H emission with FWHMs km s-1, along with broad Paschen and He I lines. All of them exhibit blueshifted He I 10830 absorbers; two of them show H absorbers, and one even shows a redshifted H absorber without an H counterpart. None of the objects shows significant optical or IR variability. Two were targeted in hard X-rays but were either not detected or detected with low statistical significance.
The Balmer and Paschen decrements, covering from H to H12 and from Pa to Pa12 (or Pa9 for one object), indicate negligible dust attenuation in the NLRs. The broad H/H ratios are in all three objects, higher than those observed in most local AGNs, indicating either unusually high-density gas or atypical dust properties in the BLRs.
The BPT diagram shows that the narrow emission lines are photoionized by AGNs and overlap with the region occupied by a known LRD. The NLRs are metal-poor, with 12+(O/H)7.43–8.00.
Two of the objects exhibit prominent narrow [Fe II] emission lines that are distinct from those typically seen in type-1 AGNs. Photoionization modeling suggests that the [Fe II]-emitting gas is AGN-powered, with physical conditions characterized by , , , and an average temperature of K. Based on the inferred gas density and FWHMs, we propose that the [Fe II]-emitting region lies at an intermediate scale between the BLRs and NLRs.
One of the local LRDs, J1025+1402 (The Egg) exhibits exceptionally high EWs in Na D, K I, Fe II, and Ca T absorption lines, with narrow FWHMs of about 300–400 km s-1 and no significant velocity shifts. Such strong metal absorbers are rarely seen in the ISM, galaxies, AGNs, or stars. Its Ca T EW is matched only by that of two supergiant and hypergiant stars. These extreme properties suggest a non-stellar and non-ISM origin. We interpret the Ca T absorption in The Egg as arising from a cool gas envelope around the BH, analogous to stellar atmospheres. We infer – and at the photosphere, corresponding to a density of g cm-3.
The optical-IR continua of these local LRDs exhibit pseudo-blackbody shapes. The shapes are consistent with theoretical models featuring atmospheres surrounding BHs (H. Liu et al., 2025), with characteristic temperatures of K. The WISE photometry indicates modest dust emission at temperatures of – K.
By combining the observational evidence, we propose a conceptual model for the structure of LRDs. In this scenario, BHs are embedded within a cool gas envelope that emits thermalized radiation, giving rise to the observed pseudo-blackbody continuum and metal absorption features. The Balmer and He I absorbers are attributed to optically thick outflows launched from the envelope. The overall AGN structure consists of BLRs, an intermediate region that contains [Fe II]-emitting gas, and NLRs with minimal dust attenuation. The geometry of these structures, along with the viewing angle, may account for the diverse observed features across different objects.
Local LRDs offer a valuable opportunity to investigate the physical nature of this population. The conceptual framework presented in §7 provides one possible scenario to explain the observed features. However, open questions remain. Do LRDs, and the scenario illustrated in Figure 14, represent a specific evolutionary phase of SMBHs? Are they connected to conventional type-1 and type-2 AGNs? How do the physical conditions that give rise to LRDs evolve with redshift, and why are they so rare in the local Universe? How do the BH and host galaxy properties of LRDs change over cosmic time? While answering these questions is beyond the scope of this paper, they are critical for understanding the broader context of BHs and merit future investigation.
Ongoing wide-field surveys, such as DESI and Euclid, along with upcoming missions like the Roman Space Telescope, will open new windows for identifying LRDs at low redshift. These efforts will enable studies of their cosmic evolution and contributions to the history of SMBH growth. Large statistical samples will be essential for establishing their demographics. Future observations with high-resolution facilities such as HST and JWST will provide critical insights to test the proposed scenarios and directly probe the properties of their host galaxies.
Appendix A HST FR782N imaging of J1047+0739
As described in §4.1, HST program GO-15617 (PI: F. Bauer) obtained high resolution ACS FR782N ramp filter images on J1047+0739. To cover the continuum, broad H, and narrow H lines, the observations were taken using the FR782N ramp filter with four configurations centered at 7474Å 7583Å 7669Å, and 7755Å, respectively. The total integration time of each configuration is s. In our analysis, we focus on the narrow-band images centered at 7474Å and 7669Å. The 7474Å narrow-band image serves as the continuum map at rest-frame 6397Å. We subtract it from the 7669Å narrow-band image to obtain the pure H line map without continuum contribution. The H line map covers the entire narrow H component and 91% of the broad component.
We compare the profile of the continuum and H line map with that of the PSF. The PSF is constructed using a star within the HST field-of-view 16″ away from the object. As shown in the left panel of Figure A.1, the continuum map has a bright central point source and an extended asymmetric structure, which we assume to be stellar light from the LRD host galaxy. Its radial profile is consistent with a PSF profile within kpc, while showing a deviation at kpc. The H line map exhibits a point-source-like morphology with clear diffraction spikes. Its radial profile closely matches that of the PSF. It indicates that the H line is primarily contributed by the AGN.
We model the morphology of the continuum map using galfit (C. Y. Peng et al., 2002). We assume a PSF and four Sérsic components. The right panel of Figure A.1 presents the best-fit model and the PSF-subtracted components, which represent the host galaxy. The host galaxy has compact, yet clumpy morphology, with a half-light radius of 530 pc. The luminosity of the central PSF is 5.5 times that of the host galaxy, indicating that the AGN contributes about 85% of the total continuum light in the optical.
We note that the compact host galaxy of J1047+0739 detected by HST at will be very challenging to resolve and detect for an LRD at high redshift, even at JWST resolution. This highlights the need for spatially resolved studies of local LRDs.
Appendix B Emission line fitting
For strong emission lines with visible broad components (e.g., H, Pa), we model each line with a single narrow Gaussian plus one or more broad Gaussian or Lorentzian profiles to capture the composite broad features. A linear polynomial is added to account for the local continuum. Specifically, broad Paschen lines are modeled with Lorentzian profiles, leveraging the high-resolution FIRE spectra to resolve line wings, while broad H components are depicted by two to three Gaussian profiles. The [N II] 6548, 6583 doublet is fit simultaneously with H. We model them with two narrow Gaussians, fixing their FWHMs to match the narrow H component and setting their flux ratio to 3. In J1047+0739, the Paschen lines and adjacent IR He I lines are modeled simultaneously, with the He I emission represented by broad and narrow Gaussian components. Note that for each line profile fit, the parameters of the narrow Gaussian component are treated independently, as the spectral resolution varies across wavelengths and instruments. For example, [O III]5007 in the MODS spectra is unresolved, with its observed FWHM determined by the wavelength-dependent line spread function. Therefore, we do not use its fit to constrain the narrow components of H or Pa.
For the H emission lines of The Egg and J1022+0841, whether from MODS or Binospec spectra, we simultaneously fit the absorption and emission profiles. Given the higher resolution of the Binospec spectra, we rely on them for more precise absorber measurements. As shown in Figure B.1, a Gaussian absorption component (EW Å) is applied on top of the composite broad H profile (i.e., the sum of the two broad Gaussians). For J1022+0841, we test two different fitting models: one assuming a single Gaussian absorber superimposed on the broad emission lines, and the other using two Gaussian absorbers. As shown in Figure B.1, the two-Gaussian model provides a better fit, more accurately capturing the profile wing.
Since the He I 10830 absorption profiles are complex, we model the intrinsic emission by masking the absorption-affected regions. The remaining line profile wings, unaffected by absorption, are used to constrain and recover the symmetric emission line profiles. The emission is modeled with one or two broad Gaussian components plus a narrow component. For H line of The Egg, we also mask the absorbed region and fit the intrinsic emission profile.
The normalized H absorption profiles shown in Figure 6 are derived by dividing the absorbed broad H emission by the intrinsic broad H line profile. For H in The Egg and He I 10830 in all three objects, the profiles are normalized by the total unabsorbed profiles, including both the broad emission lines and continua. Normalizing H using only the broad emission line is tested but results in an absorption peak below zero. We then model the normalized H absorber with a single Gaussian and the normalized He I 10830 absorber with multiple Gaussians. Table 4 summarizes the best-fit parameters for each absorber.
Name | H | H | He I 10830 | ||||||
FWHM | EW | FWHM | EW | FWHM | EW | ||||
(km s-1) | (km s-1) | (Å) | (km s-1) | (km s-1) | (Å) | (km s-1) | (km s-1) | (Å) | |
The Egg | 73 | ||||||||
(J1025+1402) | |||||||||
total | |||||||||
J1047+0739 | – | – | – | – | – | – | |||
total | |||||||||
J1022+0841 | – | – | – | (two Gaussian) | |||||
total | |||||||||
(one Gaussian) | |||||||||
Appendix C [Fe II] modeling
To model AGN photoionization scenarios for [Fe II] emission, we adopt AGN SEDs from C. Jin et al. (2012). While the chosen SED corresponds to an Eddington ratio of 0.17, our cloudy modeling is intended to be qualitative and is not highly sensitive to this assumed value. How the Eddington ratio impacts the AGN SED, and consequently the resulting [Fe II] spectrum, is beyond the scope of this paper. Our cloudy models span a grid of hydrogen densities from to (0.5 dex step), column densities from to (0.5 dex step), ionization parameters from to (1 dex step), and metallicities from to (1 dex step). We run a suite of cloudy models over this parameter space, assuming a plane-parallel geometry, and obtain the predicted [Fe II] line spectra.
To determine the best-fit models, we compare the observed emission line fluxes with those predicted by the modeled [Fe II] spectrum. Specifically, we use the fluxes of [Fe II] 4235, 4244, 4278, 4287, 4416, 5158, 5262, 7155, and 7172. Additionally, since [Fe II] 5169 is highly sensitive to the gas conditions but is not detected, we include its 3 upper limit as an extra constraint. All line fluxes, both observed and modeled, are normalized to that of [Fe II] 4244, and we compute the value for each model grid.
For comparison, we also explored models using a 5000 K blackbody as the ionizing source. The temperature is chosen based on our analysis in §6 and §7. These models fail to reproduce the observed [Fe II] line ratios, with significant discrepancies particularly in [Fe II] 4235, 4244, 4278, 4287, and 4416.
In Figure C.1, we present the best-fit [Fe II] model for J1022+0841. Its best-fit parameters are identical to those used for J1025+1402.
Appendix D The parameterization of the continua
D.1 Theoretical atmosphere model + dust templates
We begin with the best-matched theoretical models from H. Liu et al. (2025) and model the dust emission using the dust templates from J. Lyu & G. H. Rieke (2021). These dust templates, originally developed for type-1 AGNs, include four components: blackbodies at 2100 K, 1038 K, and 300 K, and a graybody at 90 K. We extrapolate the theoretical model spectra by assuming a blackbody tail in the near-IR, consistent with the model’s effective temperature. We then fit the amplitudes of the four dust components to reproduce the overall continuum and WISE photometry at wavelengths longer than 1.5 µm.
Figure 13 demonstrates that this model successfully reproduces the near-infrared photometry of The Egg. Figure D.1 shows the SEDs of J1047+0739 and J1022+0841 compared with and the models from H. Liu et al. (2025) that visually match the data. The intrinsic luminosity of the H. Liu et al. (2025) model is erg s-1 for J1047+0739, with a reference density of g cm-3 (which is comparable to the photosphere density; see H. Liu et al. (2025) for exact definition). These model parameters give an effective temperature of K. For J1022+0841, the model luminosity is erg s-1 with g cm-3, and the resulting K. We note that the SED of J1047+0739 is less well reproduced, mainly due to the non-negligible galaxy light. The HST narrow-band imaging (see Appendix A) reveals that, although the AGN dominates the continuum, the host galaxy contributes at least 18% of the continuum light around the H wavelengths. The contribution from the host galaxy is likely minor in The Egg and J1022+0841, both of which exhibit more prominent V-shaped SEDs with redder optical continua. This suggests stronger AGN dominance in the optical continua. Nevertheless, the comparison between the model and observed spectra aims to demonstrate the plausibility of the theoretical models rather than to constrain the physical properties of these objects.
As shown in Figure D.1, the rest-frame near-IR emission of J1047+0739 and J1022+0841 can be described by dust components with temperatures ranging from 2100 K to 90 K. However, the inferred amplitudes are highly sensitive to the assumed optical-to-IR continuum shape and the choice of dust templates. We discuss alternative approaches and associated degeneracies below.
D.2 An empirical parameterization of the continuum
We attempt to empirically parameterize the optical-IR continuum shapes of local LRDs with a pseudo-blackbody spectrum. This parameterization is purely descriptive, intended to capture the observed continuum shape but not to provide a physical interpretation. We adopt the functional form of a modified Calzetti law (S. Noll et al., 2009), but it should not be interpreted as indicating any particular amount of dust attenuation. The parameterization can be expressed as:
where the free parameters are , , , , , , and . and are the amplitude and power-law slope of the UV continuum. It represents a distinct component, which could be either due to the AGN light or the host galaxy. denotes the form of the modified Calzetti law (S. Noll et al., 2009), where represents a generalized attenuation curve (not necessarily tied to physical dust properties), and is the power-law modification parameter, normalized at 5500Å. For J1047+0739, we introduce an additional free parameter to model the Balmer jump as a discontinuity in the power-law at 3645 Å. is the blackbody spectrum at temperature , and is its amplitude. For the dust emission, we adopt the dust templates from J. Lyu & G. H. Rieke (2021). and are the amplitude and blackbody or gray body spectrum of the -th dust component.
The best-fit continua of the three local LRDs are shown in Figure D.2. The corresponding best-fit parameters and component luminosities are summarized in Table 5. We do not include associated uncertainties given the descriptive nature of this parameterization. All three objects exhibit temperatures around 5000–6000 K and very negative values of . The extreme produces a very steep modification to the original blackbody. Such a negative strongly suppresses the UV-to-optical flux of the blackbody spectrum while leaving the IR spectrum largely unaffected. As discussed in §6.2, physical mechanisms, such as metal cooling and gas opacity, may contribute to this modification. In particular, the contribution of the UV power-law component to the optical continuum in J1047+0739 is non-negligible. At the H wavelength, the power-law component is only a factor of 2.3 fainter than the pseudo-blackbody. This is broadly consistent with measurements from its HST narrow-band image, which show the AGN luminosity to be 5.5 times that of the host galaxy. In the other two objects, the power-law components at optical wavelengths are orders of magnitude fainter than the pseudo-blackbody components.
Parameter | J1025+1402 (The Egg) | J1047+0739 | J1022+0841 |
( erg s-1 Å-1) | |||
( erg s-1) | |||
( erg s-1) | – | – | |
( erg s-1) | |||
( erg s-1) | |||
( erg s-1) |
In The Egg, the two most prominent dust components are the 1038 K and 300 K blackbody components, which dominate the SED at wavelengths of 3-10 µm and 10-20 µm, respectively. They drive the rising flux observed in the four WISE bands. The contributions from the remaining two dust components are minor.
D.3 The Haro 11 dust template
The inferred dust components are highly sensitive to the choice of fitting templates. In J1047+0739, the W2 photometry at a rest-frame wavelength of 4 µm is not well reproduced, as simple blackbody or graybody models fail to capture features such as PAH emission and silicate absorption. In J1022+0841, a dust component with K appears, but this primarily reflects the influence of the four dust templates adopted from J. Lyu & G. H. Rieke (2021). To explore alternative interpretations, we replace the dust component in Equation D1 with the Haro 11 dust template from J. Lyu et al. (2016) and fit the continua of the three local LRDs, as shown in Figure D.3. Notably, the dust emission in J1022+0841 is well reproduced by the Haro 11 template, suggesting that its dust may reside in the ISM and be heated by AGN-driven feedback or massive star clusters in the host galaxy. This degeneracy highlights the need for mid- to far-infrared observations to robustly constrain the dust content and its origin in these systems.

### v2

@@
 The emission line analyses suggest narrow lines originate from AGN-powered, metal-poor regions with minimal dust; broad lines come from inner regions with exceptionally high density or atypical dust properties; and [Fe II] emission arises from dense gas between broad and narrow-line regions.
-One of our objects, J1025+1402 (nicknamed The Egg), shows extremely high equivalent width Na D, K I, Fe II, and Ca II triplet absorption lines, suggesting the presence of a cool (5000 K), metal-enriched gas envelope.
+One of our objects, J1025+1402 (nicknamed The Egg), shows extremely high equivalent width Na D, K I, and Ca II triplet absorption lines, along with other potential low-ionization absorption features, suggesting the presence of a cool (5000 K), metal-enriched gas envelope.
 The optical/near-IR continua of these LRDs are also consistent with theoretical models featuring an atmosphere around black holes.
@@
 1 Introduction
-The so-called “little red dots” (LRDs) have been among the most significant discoveries of the early years of JWST operations (e.g., I.
-Labbe et al., 2025; J.
-Matthee et al., 2024; X.
-Lin et al., 2024; H.
-B.
-Akins et al., 2024).
-These sources appear compact even in JWST imaging, and the majority of them (over 70–80%) exhibit broad Balmer emission lines with FWHM km s-1, indicative of black holes (BHs) of – (e.g., J.
-E.
-Greene et al., 2024; D.
-D.
-Kocevski et al., 2024; J.
-Zhang et al., 2025).
-They display V-shaped spectral energy distributions (SEDs), characterized by a blue UV continuum and a red optical continuum (D.
-D.
-Kocevski et al., 2024; D.
-J.
-Setton et al., 2024; R.
-E.
-Hviding et al., 2025).
-Moreover, a large fraction of these V-shaped objects present Balmer absorption on top of the broad emission line, which is rarely seen in other type-1 AGNs (e.g., J.
-Matthee et al., 2024; X.
-Lin et al., 2024; B.
-Wang et al., 2025a; F.
-D’Eugenio et al., 2025).
-Debates about the nature of LRDs are ongoing, particularly concerning the origin of their V-shaped SEDs, the spectral shape inflection near the Balmer break, and the prevalence of Balmer absorption (e.g., D.
-J.
-Setton et al., 2024; X.
-Lin et al., 2024).
-Early studies have shown that if the Balmer break arises from massive host galaxies, it would challenge our current understanding of structure formation (e.g., B.
-Wang et al., 2024; I.
-Labbe et al., 2024).
-Alternatively, if these features are attributed to reddened AGNs, it is difficult to account for the lack of strong rest-frame IR detections for most high-redshift LRDs, by JWST/MIRI, Spitzer, or ALMA (e.g., P.
-G.
-Pérez-González et al., 2024; C.
-C.
-Williams et al., 2024; Y.
-Ma et al., 2024; D.
-J.
-Setton et al., 2025).
-New models have started to explore the distribution of dust (Z.
-Li et al., 2025; K.
-Chen et al., 2025), along with new constraints on the dust content in LRDs from multi-wavelength data (C.
-M.
-Casey et al., 2024, 2025).
-The unusual continuum shape of LRDs has motivated multiple theoretical models (K.
-Inayoshi et al., 2025; D.
-Kido et al., 2025).
-Recent studies have proposed that BHs embedded in dense gas with a high covering fraction can explain the presence of both the observed Balmer breaks and absorption features (K.
-Inayoshi & R.
-Maiolino, 2024).
-Such models have been applied to several high-redshift LRDs and have achieved reasonably good fits to their spectra (X.
-Ji et al., 2025; A.
-J.
-Taylor et al., 2025a; A.
-de Graaff et al., 2025; R.
-P.
-Naidu et al., 2025).
+The so-called “little red dots” (LRDs) have been among the most significant discoveries of the early years of JWST operations (e.g., Labbe2025; Matthee2024; Lin2024; Akins2024).
+These sources appear compact even in JWST imaging, and the majority of them (over 70–80%) exhibit broad Balmer emission lines with FWHM km s-1, indicative of black holes (BHs) of – (e.g., Greene2024; Kocevski2024; Zhang2025).
+They display V-shaped spectral energy distributions (SEDs), characterized by a blue UV continuum and a red optical continuum (Kocevski2024; Setton2024; Hviding2025).
+Moreover, a large fraction of these V-shaped objects presents Balmer absorption on top of the broad emission line, which is rarely seen in other type-1 AGNs (e.g., Matthee2024; Lin2024; Wang2025; DEugenio2025).
+Debates about the nature of LRDs are ongoing, particularly concerning the origin of their V-shaped SEDs, the spectral shape inflection near the Balmer break, and the prevalence of Balmer absorption (e.g., Setton2024; Lin2024).
+Early studies have shown that if the Balmer break arises from massive host galaxies, it would challenge our current understanding of structure formation (e.g., Wang2024; Labbe2024).
+Alternatively, if these features are attributed to reddened AGNs, it is difficult to account for the lack of strong rest-frame IR detections for most high-redshift LRDs, by JWST/MIRI, Spitzer, or ALMA (e.g., Perez-Gonzalez2024; Williams2024; Ma2024; Xiao2025; Setton2025).
+New models have started to explore the distribution of dust (Li2025; Chen2025), along with new constraints on the dust content in LRDs from multi-wavelength data (Casey2024; Casey2025).
+In addition to their weak infrared and sub-millimeter emission, the weak X-ray (Yue2024; Ananna2024; Sacchi2025) and radio emission (Mazzolari2024; Perger2025), further set LRDs apart from typical type-1 AGNs (Kokubo2024; Maiolino2025).
+The unusual continuum shape of LRDs has motivated multiple theoretical models (Inayoshi2025b; Kido2025).
+Recent studies have proposed that BHs embedded in dense gas with a high covering fraction can explain the presence of both the observed Balmer breaks and absorption features (Inayoshi2024).
+Such models have been applied to several high-redshift LRDs and have achieved reasonably good fits to their spectra (Ji2025; Taylor2025; deGraaff2025; Naidu2025).
 However, for sources with rest-frame IR detections, these models struggle to simultaneously reproduce the weak IR photometry and the shape of Balmer breaks.
-They typically require either an extremely high column density of hydrogen gas (e.g., ), exotic modified dust attenuation laws, or intrinsically red AGN SEDs (A.
-de Graaff et al., 2025).
-Even with these additional free parameters, challenges still remain for fully explaining the IR broadband photometry (e.g., A.
-de Graaff et al., 2025; R.
-P.
-Naidu et al., 2025; A.
-J.
-Taylor et al., 2025a).
-Currently, only a limited number of high-redshift LRDs have been detected with JWST/MIRI, while the vast majority remain either not covered by or undetected in the mid-IR (e.g., G.
-C.
-K.
-Leung et al., 2024; D.
-J.
-Setton et al., 2025; X.
-Lin et al., 2025; A.
-J.
-Taylor et al., 2025a).
-Most JWST/NIRSpec spectroscopy of LRDs at has been conducted in the prism mode, and high-resolution grating observations of several objects reveal only their brightest features.
+They typically require either an extremely high column density of hydrogen gas (e.g., ), exotic modified dust attenuation laws, or intrinsically red AGN SEDs (deGraaff2025).
+Even with these additional free parameters, challenges remain for fully explaining the IR broadband photometry (e.g., deGraaff2025; Naidu2025; Taylor2025).
+Currently, only a limited number of high-redshift LRDs have been detected with JWST/MIRI, while the vast majority remain either not covered by or undetected in the mid-IR (e.g., Leung2024; Setton2025; Lin2025; Taylor2025).
+Most JWST/NIRSpec spectroscopy of LRDs at has been conducted in the prism mode, and high- or medium-resolution grating observations of several objects reveal only their brightest features.
 In contrast, LRDs at low redshifts can be observed with a variety of facilities/instruments at multiple wavelengths, with high spectral and spatial resolution at relatively low cost, providing deep insights into the nature of LRDs as a population.
-Rest-frame IR and high-resolution grating spectra have been obtained for a small number of luminous LRDs at cosmic noon (e.g., I.
-Juodžbalis et al., 2024; B.
-Wang et al., 2025a; A.
-de Graaff et al., 2025).
-Systematic searches for LRDs at using wide-field sky surveys are now underway and have already demonstrated successful spectroscopic follow-up with ground-based facilities (R.
-Lin et al., 2025; Y.
-Ma et al., 2025; Euclid Collaboration et al., 2025).
+Rest-frame IR and high-resolution grating spectra have been obtained for a small number of luminous LRDs at cosmic noon (e.g., Juodzbalis2024; Wang2025; deGraaff2025).
+Systematic searches for LRDs at using wide-field sky surveys are now underway and have already demonstrated successful spectroscopic follow-up with ground-based facilities (RLin2025; Ma2025; LRD_EuclidCollaboration2025).
 However, photometrically selected LRDs at cosmic noon are still relatively faint ( mag), making follow-up observations challenging with current ground-based facilities.
 Identifying LRDs in the local Universe opens a valuable window into understanding the nature of this population.
-Recent observations and theoretical predictions suggest that their number density declines significantly at , dropping by 1–2 dex from to (D.
-D.
-Kocevski et al., 2024; Y.
-Ma et al., 2025; K.
-Inayoshi, 2025).
+Although the measurements of the evolution of LRD number density across cosmic time vary across different photometric and spectroscopic selections, recent observations and models show a decline in the LRD number density toward low redshift.
+(Ma2025; LRD_EuclidCollaboration2025; Loiacono2025; Inayoshi2025a).
 This implies that local LRDs are rare and are linked to physical conditions that may be more common in the early Universe.
@@
 These observations would not only be cost-effective but also may unravel details that remain challenging to obtain from their high-redshift counterparts, even with JWST.
-In this paper, we present the search and discovery of three local LRDs at =0.1–0.2 from the SDSS spectroscopic database, followed by high signal-to-noise radio ground-based spectroscopic observations.
+In this paper, we present the search and discovery of three local LRDs at =0.1–0.2 from the SDSS spectroscopic database, followed by high signal-to-noise (S/N) ratio ground-based spectroscopic observations.
 The paper is organized as follows.
@@
 Throughout this work, we adopt the AB magnitude system for all photometric measurements.
+All wavelengths and line names are given in the vacuum frame.
 All equivalent widths (EWs) are reported in the rest frame.
@@
 2 Selection and sample from SDSS
-In the literature, a wide range of selection criteria have been used to define the LRD population (e.g., K.
-N.
-Hainline et al., 2025; R.
-E.
-Hviding et al., 2025).
+In the literature, a wide range of selection criteria have been used to define the LRD population (e.g., Hainline2025; Hviding2025).
 In this paper, we adopt a narrow definition of LRDs, referring specifically to objects that simultaneously satisfy the following criteria: (1) presence of broad Balmer emission lines; (2) compact morphology in the rest-frame optical, particularly in photometric bands dominated by H emission; and (3) a V-shaped SED, characterized by a blue UV continuum slope and red optical continuum slope.
@@
 We first built a library of high-redshift LRDs with well-characterized UV-to-optical spectra.
-We adopted the compilation of high-redshift LRDs from D.
-J.
-Setton et al.
-(2024), which exhibit clear V-shaped SEDs with inflection points near the Balmer limit.
+We adopted the compilation of high-redshift LRDs from Setton2024, which exhibit clear V-shaped SEDs with inflection points near the Balmer limit.
 Their JWST/NIRSpec PRISM spectra were obtained from the DAWN JWST Archive 111https://dawn-cph.github.io/dja/.
-We then searched for local LRD candidates in the SDSS DR17222https://www.sdss4.org/dr17/ spectroscopic database (K.
-N.
-Abazajian et al., 2009; Abdurro’uf et al., 2022).
-We started with the GUVmatch333https://archive.stsci.edu/hlsp/guvcat catalog, which cross-matches GALEX far-UV (FUV) and near-UV photometry (NUV) with SDSS DR14 ugriz photometry.
-We shifted the high-redshift LRD PRISM spectra to –0.5 to serve as templates for our local LRD search.
+We then searched for local LRD candidates in the SDSS DR17222https://www.sdss4.org/dr17/ spectroscopic database (Abazajian2009; Abdurrouf2022).
+We started with the GUVmatch catalog (GUVcat_DOI), which cross-matches GALEX far-UV (FUV) and near-UV photometry (NUV) with SDSS DR14 ugriz photometry.
+We shifted the high-redshift LRD PRISM spectra to –0.5, serving as templates for our local LRD search.
 This redshift range was chosen to ensure that the H emission line falls within the SDSS spectroscopic wavelength coverage.
@@
 Then, we selected sources with a distance of less than one magnitude from any of the templates.
-Finally, we imposed a compactness criterion by requiring the SDSS-reported Petrosian radii (V.
-Petrosian, 1976) in both the g and r bands to be smaller than 1.8 arcsec (physical scale of 3.3 kpc at ).
+Finally, we imposed a compactness criterion by requiring the SDSS-reported Petrosian radii (Petrosian1976) in both the g and r bands to be smaller than 1.8 arcsec (physical scale of 3.3 kpc at ).
 This morphological criterion was relatively permissive, and variations in the threshold do not affect the selection results.
 After the photometric selection, we retrieved the SDSS spectra of the selected candidates.
-We examined their emission line properties using the SPZLINE extension, which provides automated emission-line fits from the SDSS pipeline (A.
-S.
-Bolton et al., 2012).
-We first required the EWs of [O III] 5008 to exceed 10 Å, and the line ratio [O III] 5008/[O II] 3727 (O32) to be greater than 10.
+We examined their emission line properties using the SPZLINE extension, which provides automated emission-line fits from the SDSS pipeline (Bolton2012).
+We first required the EWs of [O III] 5008 to exceed 10 Å, and the line ratio [O III] 5008/[O II] 3730 (O32) to be greater than 10.
 The first criterion effectively excludes stellar objects, quiescent galaxies, and post-starburst galaxies, which could exhibit strong Balmer breaks and are prevalent in the local Universe.
-The second criterion ensures that the selected galaxies have physical conditions similar to those of high-redshift galaxies.
+The second criterion ensures that the selected galaxies have physical conditions similar to those of high-redshift galaxies (e.g., Shapley2023; Sanders2023; Shapley2025).
 To further remove contaminants and confirm the presence of red optical continua, we inspected the continuum underlying emission lines in the SPZLINE extension.
 We required the continuum under H to be redder than that under [O III] 5008, i.e., in space.
-We also required the [O III] 5008 continuum to be redder than that under [O II] 3727.
-Although we define LRDs by the presence of broad H lines, up to this stage, the broad-line criterion has not yet been explicitly applied during the selection process above.
-Nonetheless, all targets selected based on the above criteria have already exhibited clear broad H emission in their SDSS spectra.
-We therefore obtained a sample of local sources that meet all the criteria defining LRDs from the SDSS DR17 spectroscopic database.
+We required the [O III] 5008 continuum to be redder than that under [O II] 3730.
+After obtaining the initial parent sample based on the automated pipeline products, we visually inspected the candidates to exclude objects affected by pipeline uncertainties.
+For example, we removed cases where the more precise re-measured [O II] flux remained clearly stronger than [O III] , or where emission lines had low S/N and were dominated by noise.
+This process resulted in a parent sample of 40 objects.
+Although the broad-line H criterion, which is one of the defining characteristics of LRDs, was not explicitly applied up to this stage, all selected targets already show clear broad H emission in their SDSS spectra.
+We note, however, that not all objects in the parent broad-line sample satisfy the LRD definition.
+Many sources in the parent sample show very strong [N II] or Mg II emission accompanied by relatively red optical continua, in which case the spectral shapes are more naturally explained by type-1 AGNs with evolved stellar populations.
+In the end, we select three objects as the most representative LRDs.
 2.2 Sample from SDSS
@@
 Figure 1 shows the images, photometry, and spectra of the three selected local LRDs. These objects closely match their corresponding LRD templates and exhibit unambiguous V-shaped UV-to-optical SEDs.
-In fact, the local LRDs presented in this paper were originally selected as quasar candidates (G.
-T.
-Richards et al., 2002).
+The local LRDs presented in this paper were originally selected as quasar candidates (Richards2002).
 The SDSS target selection flag for J1025+1402 and J1022+0841 is QSO_HIZ, as their colors meet the criteria for selecting high-redshift quasars, primarily at .
@@
 J1047+0739 is flagged as both QSO_CAP and QSO_HIZ: the former because its color meets the selection criteria for quasars at , and the latter because it satisfies a relaxed criterion for high-redshift quasars.
-These flags indicate that the three selected objects are outliers in both the and color spaces, making them easily misclassified during color selection.
-Our LRD selection is incomplete, since it requires that the object has an existing SDSS spectrum.
-The SDSS main galaxy survey is biased against LRDs because of their compact morphology, which is one of the reasons that they are labeled as quasar candidates.
+The misidentification indicates that the three selected objects are outliers in both the and color spaces, making them easily misclassified during color selection.
+From the SDSS spectra, J1025+1402 shows clear blueshifted H absorption, while J1022+0841 exhibits tentative but unresolved H absorption.
+Both features are confirmed by our spectroscopic follow-up with higher spectral resolution (see §4.3 and Figure 5).
+J1025+1402 and J1047+0739 were identified by Izotov2008 as metal-poor AGN in SDSS DR5, with 12 + (O/H) = 7.36 and 7.99, respectively.
+Simmonds2016 further noted that they are weak in X-rays, revealed by the non-detection of Chandra.
+Burke2021 found no significant variability in either optical broadband photometry or broad H luminosity.
+2.3 Number density of LRDs at
+Our LRD selection at is incomplete.
+First of all, it requires that the object has an existing SDSS spectrum.
+The SDSS main galaxy survey is biased against LRDs because of their compact morphology, making them less likely to be prioritized for spectroscopic follow-up as galaxies.
+Their compactness is also one of the reasons that they are labeled as quasar candidates.
+In addition, although our [O III] EW criteria effectively exclude stars, quiescent, and post-starburst galaxies at low redshifts, they may also omit potential LRD candidates with intrinsically weak [O III] emission.
+Indeed, several high-redshift LRDs exhibit low [O III] EWs, such as Abell2744-QSO1 (4.5 Å; Ji2025), MoM-BH∗-1 (3 Å; Naidu2025), and The Cliff ( Å; deGraaff2025).
 Therefore, we estimate a conservative lower limit on the number density of local LRDs at =0–0.5, based on the effective SDSS survey area of 9,376 deg2.
 The discovery of these three objects yields an estimated lower limit of approximately within this redshift range.
-Our estimated number density at =0–0.5 is more than several orders of magnitude higher than the extrapolation of the log-normal distribution fitted to the LRD population at 4–7 (K.
-Inayoshi, 2025).
+Our estimated number density at =0–0.5 is more than several orders of magnitude higher than the extrapolation of the log-normal distribution fitted to the LRD population at 4–7 (Inayoshi2025a).
 However, this large discrepancy arises in the exponentially damped tail of the assumed distribution, where such divergence is expected due to the steep functional form and limited constraints at low redshifts.
-We find that a modest increase in the dispersion (by 10–20%) of the log-normal model from K.
-Inayoshi (2025) yields a redshift evolution of the LRD abundance consistent not only with measurements at 2–3 (Y.
-Ma et al., 2025), but also with our observations at =0–0.5.
-Nonetheless, this trend suggests that the BH accretion mode giving rise to LRDs, while rare in the local Universe, is not exclusively unique to the early Universe.
-From the SDSS spectra, J1025+1402 shows clear blueshifted H absorption, while J1022+0841 exhibits tentative but unresolved H absorption.
-Both features are confirmed by our spectroscopic follow-up with higher spectral resolution (see §4.3).
-J1025+1402 and J1047+0739 were identified by Y.
-I.
-Izotov & T.
-X.
-Thuan (2008) as metal-poor AGNs in SDSS DR5, with 12 + (O/H) = 7.36 and 7.99, respectively.
-C.
-Simmonds et al.
-(2016) further noted that they are weak in X-rays, revealed by the non-detection of Chandra.
-C.
-J.
-Burke et al.
-(2021) found no significant variability in either optical broadband photometry or broad H luminosity.
+We find that a modest increase in the dispersion (by 10–20%) of the log-normal model from Inayoshi2025a yields a redshift evolution of the LRD abundance consistent not only with measurements at 2–3 (Ma2025), but also with our observations at =0–0.5.
+The evolution of LRD number density across cosmic time remains uncertain, with different studies reporting varying trends.
+Photometric selection and spectroscopic follow-ups by Ma2025 indicate a substantial decline at , with the number density decreasing by 1–2 dex from to .
+In contrast, photometric selection in LRD_EuclidCollaboration2025 suggests a milder evolution, with LRD abundance rising to before declining at lower redshift.
+Based on the discovery of three LRDs at , Loiacono2025 implies that LRD number density remains significant at cosmic noon, only 2–3 times lower than that of UV-selected quasars.
+Nevertheless, both Ma2025 and LRD_EuclidCollaboration2025 indicate a decreasing number density toward the local Universe.
+This trend, together with the discovery of the three local LRDs presented in this work, suggests that the BH accretion mode giving rise to LRDs, while rare in the local Universe, is not exclusively unique to the early Universe.
 3 Data and Observations
 3.1 Photometry
-The three targets were observed by multiple wide-field sky surveys.
-We adopt photometric data from the following surveys: FUV and NUV from GALEX DR6 (L.
-Bianchi & B.
-Shiao, 2020), ugriz Petrosian photometry from SDSS DR17 (K.
-N.
-Abazajian et al., 2009; Abdurro’uf et al., 2022), PSF photometry from Pan-STARRS DR2 (H.
-A.
-Flewelling et al., 2020), mean photometry from GAIA DR3 (Gaia Collaboration et al., 2023), photometry from Legacy Surveys DR10 (A.
-Dey et al., 2019), Petrosian photometry from UKIDSS DR11PLUS (A.
-Lawrence et al., 2007), and W1, W2, W3, W4 photometry from WISE (E.
-L.
-Wright et al., 2010).
+The three targets have been observed in multiple wide-field sky surveys.
+We compile photometric measurements from the following datasets, adopting PSF photometry where available since the targets are unresolved in the images: FUV and NUV from GALEX DR6 (Bianchi2020); ugriz PSF photometry from SDSS DR17 (Abazajian2009; Abdurrouf2022); PSF photometry from Pan-STARRS DR2 (Flewelling2020); mean photometry from GAIA DR3 (Gaia2023); model photometry from Legacy Surveys DR10 (Dey2019) where PSF models are used for the three objects; PSF photometry from UKIDSS DR11PLUS (Lawrence2007); and W1, W2, W3, W4 photometry from WISE (Wright2010).
 The GALEX photometry is taken from the GUVmatch catalog.
-The WISE photometry has been incorporated into the Legacy Survey catalog based on the unWISE images (E.
-F.
-Schlafly et al., 2019).
-All other data are retrieved from Astro Data Lab444https://datalab.noirlab.edu/ (M.
-J.
-Fitzpatrick et al., 2014; R.
-Nikutta et al., 2020).
-3.2 Spectroscopic Follow-up Observations
-We obtained high signal-to-noise-ratio (S/N) LBT/MODS and Magellan/FIRE spectroscopy to characterize the UV-to-NIR spectra of the three local LRDs.
+The WISE photometry has been incorporated into the Legacy Survey catalog based on the unWISE images (Schlafly2019).
+All other data are retrieved from Astro Data Lab333https://datalab.noirlab.edu/ (Fitzpatrick_SPIE; NIKUTTA2020100411).
+3.2 Spectroscopic Follow-up Observations and Data Reduction
+We obtained high S/N LBT/MODS and Magellan/FIRE spectroscopy to characterize the UV-to-near-IR spectra of the three local LRDs.
 We also obtained MMT/Binospec spectra of J1025+1402 and J1022+0841 to measure their high-resolution H profiles.
-The LBT/MODS (R.
-W.
-Pogge et al., 2010) is a pair of dual-channel spectrographs on the 28.4m Large Binocular Telescope, with the blue channel covering 3300–6000 Å and the red channel 5000–10000 Å.
+The LBT/MODS (Pogge2010) is a pair of dual-channel spectrographs on the 28.4 m Large Binocular Telescope, with the blue channel covering 3300–6000 Å and the red channel 5000–10,000 Å.
 We observed the three targets on February 24, 2025, using 1.0-arcsec slits in grating mode under an average seeing of 0.8 arcsec.
@@
 As a result, each target effectively has 6600 s exposures in both the blue and red channels.
-The Magellan/FIRE (R.
-A.
-Simcoe et al., 2013) is a near-IR echelle spectrograph covering 0.8–2.5 on the 6.5m Magellan Baade Telescope.
+The Magellan/FIRE (Simcoe2013) is a near-IR echelle spectrograph covering 0.8–2.5 on the 6.5 m Magellan Baade Telescope.
 We observed J1025+1402 using a 0.75-arcsec slit on December 21, 2024, reaching –6000.
@@
 The seeing for these observations was –0.7 arcsec.
-The MMT/Binospec observation of J1025+1402 was conducted in the IFU mode (D.
-Fabricant et al., 2019, 2025) on December 7, 2024, using the 600 line/mm grating.
+The MMT/Binospec observation of J1025+1402 was conducted in the IFU mode (Fabricant2019; Fabricant2025) on December 7, 2024, using the 600 line/mm grating.
 The observation consisted of s exposures and achieved a spectral resolution of .
@@
 The observation included s exposures, with a spectral resolution of .
-The LBT/MODS and Magellan/FIRE data were reduced with PypeIt (J.
-Prochaska et al., 2020).
+The LBT/MODS and Magellan/FIRE data were reduced with PypeIt (Prochaska2020).
 We performed bias subtraction, flat-fielding, wavelength calibration, sky subtraction, spectral extraction, and telluric correction.
 To correct for slit losses and calibrate the absolute flux of the LBT/MODS spectra, we first matched them to the SDSS spectra using a wavelength-dependent linear function.
-The spectra in SDSS DR17 have been processed with an updated flux calibration algorithm, yielding significant improvements at wavelengths below 6000 Å555https://www.sdss4.org/dr17/algorithms/spectrophotometry/.
+The spectra in SDSS DR17 have been processed with an updated flux calibration algorithm, yielding significant improvements at wavelengths below 6000 Å444https://www.sdss4.org/dr17/algorithms/spectrophotometry/.
 We confirmed that the SDSS spectra closely follow the shape of the corresponding photometry.
@@
 The Magellan/FIRE spectra were scaled to match the UKIDSS -band photometry using a constant factor.
-The MMT/Binospec data were reduced using the Binospec pipeline (J.
-Kansky et al., 2019).
+The MMT/Binospec data were reduced using the Binospec pipeline (Kansky2019).
 The IFU data were extracted with a rectangular aperture matching the seeing disk.
-Throughout this work, most measurements are based on MODS and FIRE spectra, while Binospec spectra are used exclusively to analyze the H absorber profiles of J1025+1402 and J1022+0841.
+The Binospec spectra of J1025+1402 and J1022+0841 were flux-calibrated by scaling to match the calibrated LBT/MODS spectra after degrading them to the MODS resolution.
 The SDSS spectra have a resolution of –2500.
@@
 For J1022+0841, the S/N increases from 0.6 to 5.5 at 4000–5000 Å, and from 2.6 to 29 at 7000–8000 Å.
-Figure 1 shows the UV-to-NIR spectra, multi-wavelength photometry, and images of the three targets.
+Figure 1 shows the UV-to-near-IR spectra, multi-wavelength photometry, and images of the three targets.
 The LBT/MODS spectra confirm the presence of the V-shape of the UV-to-optical continua.
-The declining near-IR spectra in space are consistent with the rest-frame IR photometry of bright high-redshift LRDs detected by MIRI (e.g., B.
-Wang et al., 2025a; A.
-de Graaff et al., 2025; A.
-J.
-Taylor et al., 2025a; D.
-J.
-Setton et al., 2025).
+The declining near-IR spectra in space are consistent with the rest-frame IR photometry of bright high-redshift LRDs detected by MIRI (e.g., Wang2025; deGraaff2025; Taylor2025; Setton2025).
 4 Evidence as local LRDs
-Basic property
-Name | RA | DEC | (mag) | 12 + | ||||
-( erg s-1) | ( erg s-1) | ( erg s-1) | ||||||
-J102530.29+140207.3 | 156.37622 | 14.03586 | 0.1007 | 19.003 | 7.47 0.30 | 4.95 | 7.550.03 | |
-J104755.92+073951.2 | 161.98302 | 7.66423 | 0.1682 | 18.846 | 18.5 0.53 | 8.000.02 | ||
-J102208.52+084156.1 | 155.53552 | 8.69892 | 0.2227 | 20.011 | 17.4 0.31 | – | 7.430.03 |
-H emission-line and BH properties
-Name | FWHM | FWHMHα,narrow | |||
-( erg s-1) | ( erg s-1) | (km s-1) | (km s-1) | ||
-J1025+1402 | 8.68 0.13 | 1.33 0.02 | 983 10 | unresolved | |
-J1047+0739 | 27.51 0.58 | 18.86 0.15 | 949 25 | 47 3 | 6.50 0.09 |
-J1022+0841 | 48.10 0.49 | 3.43 1.10 | 1022 18 | 6.70 0.09 | |
+| Name | J102530.29+140207.3 | J104755.92+073951.2 | J102208.52+084156.1 |
+| RA | 156.37622 | 161.98302 | 155.53552 |
+| DEC | 14.03586 | 7.66423 | 8.69892 |
+| 0.1007 | 0.1682 | 0.2227 | |
+| (mag) | 19.003 | 18.846 | 20.011 |
+| ( erg s-1) | 7.57 0.31 | 18.67 0.53 | 17.78 0.31 |
+| ( erg s-1) | 4.95 | 29.930.01 | 17.72 |
+| ( erg s-1) | – | ||
+| 12 + | 7.58 0.03 | 8.01 0.02 | 7.43 0.03 |
+| (kpc) | 0.568 | ||
+| Inflection point (Å) |
+| Name | FWHM | FWHMHα,narrow | b | b | |||
+| ( erg s-1) | (km s-1) | ( erg s-1) | (km s-1) | () | (erg s-1) | ||
+| J1025+1402 | |||||||
+| J1047+0739 | |||||||
+| J1022+0841 |
 In this section, we present the observed characteristics of three local objects that closely resemble high-redshift LRDs.
-The section is organized as follows: morphology and continuum luminosities are discussed in §4.1; broad emission lines and BH masses in §4.2; Balmer and He I absorption features in §4.3; variability properties in §4.4; X-ray and radio properties in §4.5; and a summary of the observational evidence as LRDs is provided in §4.6.
-We present their zoomed-in spectra in Figures 2, 3, and 4, and summarize the derived properties in Table 1.
-4.1 Morphology and Continuum Luminosity
+The section is organized as follows: morphology and continuum luminosities are discussed in §4.1; broad H emission line and BH properties in §4.2; Balmer and He I absorption features in §4.3; variability properties in §4.4; X-ray and radio properties in §4.5; and a summary of the observational evidence as LRDs is provided in §4.6.
+We present their zoomed-in spectra in Figures 2, 3, and 4, and summarize their properties in Table 1 and 2.
+4.1 Morphology and Continuum
 All three objects exhibit very compact morphologies across the optical bands.
 They are not spatially resolved in LBT/MODS direct imaging with a seeing of 0.9″.
-This implies that the diameter of J1025+1402 is comparable to or smaller than 1.7 pkpc, while those of J1047+0739 and J1022+0841 are constrained to be smaller than 2.5 pkpc and 2.3 pkpc, respectively.
+This implies that the diameter of J1025+1402 is comparable to or smaller than 1.7 pkpc, while those of J1047+0739 and J1022+0841 are constrained to be smaller than 2.6 pkpc and 3.2 pkpc, respectively.
 J1047+0739 was observed with the HST/ACS FR782N ramp filter narrow-band imaging that covers both the H line and the adjacent continuum (GO-15617; PI: F.
 Bauer), achieving a spatial resolution of 206 pc.
-The continuum map reveals a bright central point source contributing about 85% of the light, alongside a compact yet clumpy host galaxy with a half-light radius of 530 pc.
+The continuum map reveals a bright central point source contributing about 85% of the light, alongside a compact yet clumpy host galaxy with a half-light radius of 568 pc (see Appendix A).
 The H line map is dominated by the central point source with minimal host contribution.
@@
 All three objects exhibit continuum shapes characterized by pronounced V-shapes and declining IR slopes.
+We determine the inflection points by fitting a two-component power law to a 300 Å segment of the continuum centered on the minimum within the rest-frame 3000–4500 Å range.
 The inflections occur near the Balmer limit.
-The continuum luminosity at rest-frame 5100Å () is around erg s-1.
-J1047+0739 exhibits a potential Balmer jump, presumably due to the presence of a host galaxy.
+We measure the UV continuum slopes at rest-frame 2500 Å, , by fitting a power law to the GALEX/NUV photometry and MODS spectra below 3000 Å.
+All the three objects have .
+The optical continuum slopes () are measured from the MODS spectra around rest-frame 4500 Å, after masking emission lines and potential absorption features.
+J1025+1402 and J1022+0841 exhibit steep , while J1047+0739 shows a flatter, yet still red, optical continuum slope ().
+J1047+0739 also exhibits a potential Balmer jump, presumably due to the presence of a host galaxy.
+The continuum luminosity of the three objects at rest-frame 5100 Å () is around erg s-1.
 For all three objects, the WISE photometry at rest-frame 3–20 µm reveals the presence of weak dust emission, at a wavelength range beyond what is possible with JWST observations of high-redshift LRDs.
 We integrate the continua from rest-frame 2000 Å to 20 µm to estimate the total continuum luminosity over this wavelength range, yielding – erg s-1 for all three objects.
-4.2 Broad Emission Lines and BH masses
-The three objects show clear broad profiles in H, H, Pa to Pa, and several He I lines (e.g., He I 7065, 10830).
-J1025+1402 and J1022+0841 also exhibit broad infrared O I emission lines, consistent with the line broadening originating from gas motions in the BLRs around BHs.
-The methodology of the line measurements is detailed in Appendix B.
-Specifically for the H lines, we model the narrow component using a single narrow Gaussian.
-The broad components are fitted with two to three broad Gaussians.
-The [N II] 6548, 6583 doublet is modeled with two narrow Gaussians sharing the same FWHM as the narrow H component and fixed line ratios of 3.
-For J1025+1402 and J1022+0841, additional Gaussian absorbers are applied atop the broad emission lines.
-The measurements from the MODS spectra are summarized in Table 1.
-We also report the narrow H FWHMs of J1025+1402 and J1022+0841 from the Binospec spectra, taking advantage of its higher spectral resolution.
-All three objects exhibit broad H emission with luminosities around erg s-1, and their individual broad components exhibit FWHMs ranging from km s-1 to 2000 km s-1 , with the composite FWHMs around 1000 km s-1.
-We estimate their BH masses using the H luminosity (), the composite broad-line H FWHM, and the empirical relation from J.
-E.
-Greene & L.
-C.
-Ho (2005).
+4.2 Broad H Emission Lines and BH properties
+The three objects show clear broad profiles in H, H, Pa to Pa, and several He I lines (e.g., He I 7067, 10833).
+J1025+1402 and J1022+0841 also exhibit broad infrared O I 1.29 µm emission lines, consistent with the line broadening originating from gas motions in the BLRs around BHs.
+The methodology for the line measurements is described in Appendix B.
+For the H lines, the narrow component is modeled with a single Gaussian, while the broad component is fitted with two to three broad Gaussians.
+The [N II] 6550, 6585 doublet is modeled with two narrow Gaussians sharing the same FWHM as the narrow H component and a fixed line ratio of 3.
+For J1025+1402 and J1022+0841, we adopt the MMT/Binospec H profile measurements as fiducial and confirm their consistency with the MODS spectra.
+The resulting H measurements are summarized in Table 2.
+The effective FWHM of the broad H profiles (FWHM), used to estimate BH masses, is derived from the summed profile of all broad components.
+All three objects exhibit broad H emission with luminosities around erg s-1, and their individual broad components exhibit FWHMs ranging from km s-1 to 2000 km s-1 , with the effective FWHMs around 1000 km s-1.
+We estimate their BH masses using the broad H luminosity (), the effective broad H FWHM (FWHM), and the empirical relation from Reines2015.
 We find BH masses around –.
-This scaling relation for BH mass estimates is widely used in high-redshift LRD studies (e.g., J.
-Matthee et al., 2024; X.
-Lin et al., 2024, 2025; A.
-J.
-Taylor et al., 2025b).
-However, this estimator might not be fully applicable to LRDs, if their physical nature is different from classical AGNs, particularly if their line profiles and continua arise from distinct physical mechanisms (see §7).
-The measured values in these objects are 3–6 times higher than the predicted by the – relation in J.
-E.
-Greene & L.
-C.
-Ho (2005), where refers to the total (broad + narrow) H luminosity from AGNs.
+This scaling relation for BH mass estimates is widely used in high-redshift LRD studies (e.g., Matthee2024; Lin2024; Lin2025; Taylor2025a).
+If estimated using the Greene2005 relation, the values would be 0.15–0.33 dex lower.
+However, these estimators might not be fully applicable to LRDs, if their physical nature is different from classical AGNs, particularly if their line profiles and continua arise from distinct physical mechanisms (see §7).
+The measured values in these objects are 3–7 times higher than the predicted by the – relation in Greene2005, where refers to the total (broad + narrow) H luminosity from AGNs.
+If using and FWHMHα,broad to estimate following the relations in Greene2005, the resulting values are 0.5–0.7 dex lower than those listed in Table 2.
 It remains unclear whether or reliably traces the BLR size in these systems.
+On the other hand, if the broadening of H lines is partly due to electron scattering in dense ionized gas rather than purely from BLR gas motion, the derived could be overestimated by 1–2 dex (Naidu2025; Rusakov2025).
+However, whether this scenario applies to the general LRD population remains uncertain (Juodzbalis2025; Brazzini2025).
+We note that no dust attenuation correction has been applied in the estimates above.
+The intrinsic Balmer decrements in BLRs remain uncertain because of the complex radiative transfer effects in high-density gas clouds.
+The broad H/H line ratios in the three objects are 8-34 (see §5.1 for a more detailed analysis).
+Assuming Case B recombination, although it may not hold in BLRs, such large ratios correspond to under an SMC attenuation curve (Pei1992).
+If this attenuation correction is applied, the intrinsic H luminosity could be a factor of higher than the directly measured values, implying that the BH masses would increase by dex.
+However, if energy balance holds, such a high level of dust attenuation is in tension with the weak IR emission observed by WISE.
+For reference and comparison with literature LRDs, we also list bolometric luminosities derived solely from H following Stern2012, along with the corresponding Eddington ratios, in Table 2 ( and ).
+These values are generally consistent with high-redshift LRDs, which span a wide range of from to , and from to (e.g., Maiolino2024; Greene2024; Lin2024; Juodzbalis2025).
+We emphasize, however, that these -derived estimates may not accurately reflect the intrinsic bolometric luminosities and Eddington ratios due to the uncertain physical processes in such systems and the systematics discussed above (Greene2025).
 4.3 Balmer and He I Absorption
 Figure 5 presents the line profiles of H, H, He I, and Pa for the three local LRDs.
-All three objects exhibit absorption lines in their Balmer series or He I 10830 lines.
-For H, we model the absorber as Gaussian components superimposed on the intrinsic broad H emission.
-We then simultaneously fit the absorbers, the narrow H, broad H, and the [N II] 6548, 6583 doublet.
-For H and He I 10830, absorbed regions are masked, and the intrinsic emission line profile is modeled using a narrow Gaussian combined with multiple broad Gaussians.
-The detailed fitting methodology and complete absorption measurement information are described in Appendix B.
+All three objects exhibit absorption lines in their Balmer series and/or He I 10833 lines.
+For H of J1025+1402 and J1022+0841, we model the absorber as Gaussian components superimposed on the intrinsic broad H emission.
+We then simultaneously fit the absorbers, the narrow H, broad H, and the [N II] 6550, 6585 doublet.
+For H of J1025+1402 and He I 10833 of all three objects, absorbed regions are masked, and the intrinsic emission line profile is modeled using a narrow Gaussian combined with multiple broad Gaussians.
+The detailed fitting methodology is described in Appendix B.
+The measured absorption properties are presented in Table 3.
 J1025+1402 exhibits a redshifted H absorption feature superimposed on the broad H emission and continuum, with km s-1 and Å.
-A similar redshifted H absorber is also seen in the triply imaged LRD in the Abell 2744 field (X.
-Ji et al., 2025).
+A similar redshifted H absorber is also seen in the triply imaged LRD in the Abell 2744 field (Ji2025).
 J1025+1402 and J1047+0739 both have H absorption, superimposed on the broad H emission.
@@
 Their EWs are Å and Å, respectively, yielding a total EW of Å.
-An alternative solution is one Gaussian absorber with km s-1 and EW of Å (see Appendix B).
+An alternative solution is one Gaussian absorber with km s-1 and EW of Å, but we find that this solution leaves more residuals as commented in Appendix B.
 J1047+0739 does not show clear Balmer absorption.
@@
 This suggests that the two lines might have distinct origins.
-All three targets show prominent blueshifted He I 10830 absorption.
-The He I 10830 absorber of J1025+1402 and J1047+0739 has multiple components with distinct kinematics (Figure 6).
-The line profile of the J1025+1402 He I 10830 absorber, as shown in Figure 5, can be described using three Gaussian components, blueshifted with , and km s-1, respectively.
+For reference, we also present the Pa line profile in Figure 5.
+No absorption features associated with Pa are detected, consistent with the findings of Juodzbalis2024; Juodzbalis2025.
+This suggests that the Balmer absorbers originate from the neutral hydrogen elevated to the state but not significantly populated at .
+All three targets show prominent blueshifted He I 10833 absorption.
+The He I 10833 absorber of J1025+1402 and J1047+0739 has multiple components with distinct kinematics (Figure 6).
+The line profile of the J1025+1402 He I 10833 absorber, as shown in Figure 5, can be described using three Gaussian components, blueshifted with , and km s-1, respectively.
 This profile may alternatively result from a mixture of saturated He I absorbers.
-The line profile of the J1047+0739 He I 10830 absorber can be described using five Gaussian components with , and km s-1, respectively.
-For J1022+0841, the He I 10830 absorber is well described by a single Gaussian component with km s-1.
+The line profile of the J1047+0739 He I 10833 absorber can be described using five Gaussian components with , and km s-1, respectively.
+For J1022+0841, the He I 10833 absorber is well described by a single Gaussian component with km s-1.
 They highlight the prevalence of He I absorption in local LRDs.
-As shown in Figure 6, for J1025+1402 and J1022+0841, the velocities of the He I and H absorbers are not well aligned, indicating complex outflow kinematics.
-For reference, we also show the Pa line profile in Figure 5.
-No absorption features are associated with Pa.
-This suggests that the Balmer absorbers originate from hydrogen atoms with a population elevated to the energy level, but not to the level.
+The He I line is a resonant transition from the metastable 23S state to the 23P state.
+The 23S state can be populated either through collisional excitation in high-density gas or via recombination from He+.
+It remains unclear whether the He I absorption in LRDs originates in the same dense clouds responsible for the Balmer absorption (Juodzbalis2024), where neutral He is collisionally excited into the 23S state, or instead in lower-density partially ionized outflows (Wang2025), where He is first ionized to He+ and then recombines into the 23S state.
+Although strong He I recombination lines (4473, 5877, 7067) with 23P as the lower state are detected in all three objects, the origin remains challenging to robustly constrain, as the emitting and absorbing gas clouds may not spatially co-exist.
+The complex kinematics of He I absorbers (Figure 6) may also be attributed to mixed origins.
+| Name | H | H | He I 10833 | ||||||
+| FWHM | EW | FWHM | EW | FWHM | EW | ||||
+| (km s-1) | (km s-1) | (Å) | (km s-1) | (km s-1) | (Å) | (km s-1) | (km s-1) | (Å) | |
+| J1025+1402 | 73 | ||||||||
+| total | |||||||||
+| J1047+0739 | No absorber | No absorber | |||||||
+| total | |||||||||
+| J1022+0841 | No absorber | ||||||||
+| total | |||||||||
 4.4 Variability
@@
 For the optical bands, we retrieve multi-epoch PSF-modeled photometry from the Pan-STARRS DR2 detection catalog.
-Over 1500 days (1300 days in the rest frame), we find no significant variability, with observed variation of 0.1–0.2 mag consistent with the scatter in non-variable sources (T.
-Simm et al., 2015).
-In the WISE W1 and W2 bands, we find variations of 0.25–0.5 mag over 5000 days (4000 days in the rest frame), also consistent with the typical scatter expected for non-variable sources.
-We thus conclude that no significant variability is detected in either the rest-frame optical or infrared bands, consistent with the lack of significant variability seen in high-redshift LRDs (M.
-Kokubo & Y.
-Harikane, 2024; Z.
-Zhang et al., 2025; X.
-Lin et al., 2025; W.
-L.
-Tee et al., 2025).
-C.
-J.
-Burke et al.
-(2021) reported no significant variability in the broad H luminosity of J1025+1402 and J1047+0739 over a continuous 15-year period.
-The H luminosity range of J1025+1402 in C.
-J.
-Burke et al.
-(2021) is approximately – erg s-1, and for J1022+0841 it is – erg s-1, both well consistent with our measurements.
-They also presented V-band light curves from the Catalina Real-Time Transient Survey (CRTS; A.
-J.
-Drake et al., 2009) and -band light curves from the Zwicky Transient Facility (ZTF; F.
-J.
-Masci et al., 2019) for both objects.
-Although the limited quality of the light curves precluded definitive conclusions, they found potential low-level variability consistent with the stochastic damped random walk model of AGN variability (B.
-C.
-Kelly et al., 2009; C.
-L.
-MacLeod et al., 2010).
+Over 1500 days (1300 days in the rest frame), we find no significant variability, with observed variation of 0.1–0.2 mag consistent with systematic uncertainties as well as the scatter in non-variable sources (Simm2015).
+Following Lyu2019, we construct the WISE W1 (3.4 µm) and W2 (4.6 µm) band light curves from the WISE and NEOWISE (Mainzer2011) missions.
+In the W1 and W2 bands, we find variations of 0.25–0.5 mag over 5000 days (4000 days in the rest frame), also consistent with the typical scatter expected for non-variable sources.
+We thus conclude that no significant variability is detected in either the rest-frame optical or infrared bands, consistent with the lack of significant variability seen in high-redshift LRDs (Kokubo2024; ZZhang2025; Lin2025; Tee2025).
+Burke2021 reported no significant variability in the broad H luminosity of J1025+1402 and J1047+0739 over a continuous 15-year period.
+The H luminosity range of J1025+1402 in Burke2021 is approximately – erg s-1, and for J1047+0739 it is – erg s-1, both well consistent with our measurements.
+They also presented V-band light curves from the Catalina Real-Time Transient Survey (CRTS; Drake2009) and -band light curves from the Zwicky Transient Facility (ZTF; Masci2019) for both objects.
+Although the limited quality of the light curves precluded definitive conclusions, they found potential low-level variability consistent with the stochastic damped random walk model of AGN variability (Kelly2009; MacLeod2010).
 4.5 X-ray and Radio
-C.
-Simmonds et al.
-(2016) reported net counts of Chandra for J1025+1402 in a net exposure time of 4937 s, placing an upper limit666Assuming a power law spectrum absorbed by Galactic .
-of .
+Simmonds2016 reported net counts for J1025+1402 in the 0.5–7.0 keV band with Chandra, based on a net exposure time of 4937 s, placing an upper limit555Assuming a power law spectrum absorbed by Galactic .
+on the apparent X-ray luminosity of .
 For J1047+0739, they reported net counts in 4782 s.
-We further examined the Chandra Source Catalog Release 2.1 (I.
-N.
-Evans et al., 2024), which includes more recent observations of J1047+0739, and found net counts in the 0.5–7 keV band over a 53 ks exposure.
-This corresponds to an X-ray luminosity of , although the detection is only significant at .
+We further examined the Chandra Source Catalog Release 2.1 (Evans2024), which includes more recent observations of J1047+0739, and found net counts in the 0.5–7 keV band over a 53 ks exposure.
+This corresponds to an apparent X-ray luminosity of , although the detection is only significant at 666We use WebPIMMS to convert counts into fluxes.
+.
 For J1022+0841, no Chandra observations are currently available.
-The comparison of the X-ray emission of J1025+1402 and J1047+0739 with that of typical type-1 AGNs is detailed in C.
-Simmonds et al.
-(2016).
-The upper limits or X-ray detections suggest that J1025+1402 and J1047+0739 are X-ray weak, exhibiting ratios at least 1–2 dex below those of typical type-1 AGNs with similar broad H luminosities (J.
-E.
-Greene & L.
-C.
-Ho, 2004; F.
-Panessa et al., 2006; L.-B.
-Desroches et al., 2009).
-This X-ray weakness is consistent with that seen in high-redshift LRDs (M.
-Yue et al., 2024; T.
-T.
-Ananna et al., 2024).
-We examine the radio data from the VLA Sky Survey (M.
-Lacy et al., 2020), the TIFR GMRT Sky Survey (H.
-T.
-Intema et al., 2017), and the VLA FIRST Survey (R.
-H.
-Becker et al., 1995).
+The comparison of the X-ray emission of J1025+1402 and J1047+0739 with that of typical type-1 AGNs is detailed in Simmonds2016.
+The upper limits or X-ray detections suggest that J1025+1402 and J1047+0739 are X-ray weak, exhibiting ratios at least 1–2 dex below those of typical type-1 AGNs with similar broad H luminosities (Greene2004; Panessa2006; Desroches2009).
+This X-ray weakness is consistent with that seen in high-redshift LRDs (Yue2024; Ananna2024).
+We examine the radio data from the VLA Sky Survey (Lacy2020), the TIFR GMRT Sky Survey (Intema2017), and the VLA FIRST Survey (Becker1995).
 We find no radio detections at 153 MHz, 1.4 GHz, and 3 GHz for any of the three sources.
 The corresponding upper limits for J1025+1402, J1047+0739, and J1022+0841 are: 13.8, 12.5, and 15.1 mJy at 153 MHz; 0.7, 0.7, and 0.8 mJy at 1.4 GHz; and 0.4, 0.5, and 0.4 mJy at 3 GHz, respectively.
-The upper limits of radio-loudness (, M.
-Sikora et al.
-2007) for the three objects are estimated to be 8, 12, and 20, respectively.
+The upper limits of radio-loudness (, Sikora2007) for the three objects are estimated to be 8, 12, and 20, respectively.
 The current radio data are not deep enough to determine whether these sources are radio-quiet ().
 4.6 Summary of Evidences as LRDs
-As discussed above, the three objects at – satisfy all the defining criteria for LRDs.
+As discussed in this section and in the previous ones, the three objects at – satisfy all the defining criteria for LRDs.
 We summarize the similarity between the observed properties of the three local sources to high-redshift LRDs as follows:
@@
 V-shaped continuum (blue UV continuum slope, red optical continuum slope) and decreasing IR continuum at µm;
-broad Balmer and Paschen emission lines, consistent with – BHs;
-weak X-ray emission;
+broad Balmer emission lines, consistent with – BHs;
+weak X-ray emission; no radio detection in the current data;
 no significant (or low-level) variability;
-high occurrence rate of Balmer and He I 10830 absorption.
-Their broad H luminosities, erg s-1, fall within the range observed in high-redshift LRDs.
+high occurrence rate of Balmer and He I 10833 absorption.
+Their broad H luminosities, erg s-1, fall within the range observed in high-redshift LRDs (e.g., Matthee2024; Maiolino2024; Lin2024; Greene2024; Zhang2025; Hviding2025).
 If redshifted to , their expected apparent magnitudes in JWST/NIRCam F444W would be 25.7–26.8 mag.
-These values lie within the magnitude range observed for the diverse LRDs discovered at (e.g., J.
-E.
-Greene et al., 2024; D.
-D.
-Kocevski et al., 2024; H.
-B.
-Akins et al., 2024; D.
-D.
-Kocevski et al., 2025).
+These values lie within the magnitude range observed for the diverse LRDs discovered at (e.g., Greene2024; Kocevski2024; Akins2024; Kocevski2025).
 5 Emission Line Analysis
@@
 For high-order hydrogen lines with insufficient S/N to detect reliable broad components, we fit them with a single Gaussian representing the narrow component, assuming that the broad components are buried in the noise.
-All measurements are based on the MODS spectra for consistency.
-As shown in Figure 7, the narrow Balmer lines from H to H12 in J1025+1402 and J1022+0841 closely match the intrinsic Case B recombination predictions (, ).
-The narrow Balmer lines in J1047+0739 are consistent with the SMC dust attenuation law (Y.
-C.
-Pei, 1992) with modest attenuation of .
-The exception is H and H8, which are blended with [Ne III] and He I 3889, respectively.
+As shown in Figure 7, the narrow Balmer lines from H to H12 in J1025+1402 and J1022+0841 closely match the intrinsic Case B recombination predictions (, ; Brocklehurst1971; Hummer1987; Osterbrock2006).
+The narrow Balmer lines in J1047+0739 are consistent with the SMC dust attenuation law (Pei1992) with modest attenuation of .
+The exception is H and H8, which are blended with [Ne III] 3969 and He I 3890, respectively.
 The agreement indicates that the narrow hydrogen lines originate from regions with minimal dust attenuation.
@@
 In contrast, the He I emission is particularly strong in J1047+0739, allowing both the He I and Paschen lines to be reliably fitted.
-The broad H/H line ratios are in all three objects.
+The broad H/H line ratios are for J1025+1402, for J1047+0739, and for J1022+0841.
 Such large decrements may be attributed to either radiative transfer effects in high-density gas (e.g., collisional excitation and de-excitation, resonance scattering, etc.) or substantial dust attenuation.
-Notably, the measured decrements, both in the broad and narrow components, are well consistent with those observed in high-redshift LRDs (M.
-Brooks et al., 2025).
+Notably, the measured decrements of the broad Balmer lines are well consistent with those observed in high-redshift LRDs (Brooks2025; D'Eugenio2025_z7p04).
 These values are comparable to those expected for the SMC attenuation law with under Case B conditions.
-However, one caveat in attributing the broad-line decrements to dust is the survival of dust grains in the BLRs, which likely lies within the sublimation zone (e.g., P.
-G.
-Martin & G.
-J.
-Ferland, 1980; M.
-Kishimoto et al., 2007; C.
-M.
-Gaskell & P.
-Z.
-Harrington, 2018; A.
-Baskin & A.
-Laor, 2018).
-For comparison, type-1 AGNs in the SDSS typically show broad H/H ratios around 3.7, with the most extreme cases reaching 5 (D.
-Ilić et al., 2012).
-The much larger broad Balmer decrements observed in LRDs suggest that their BLRs may have extreme conditions with unusually dense gas or atypical dust properties, which differ substantially from those in typical type-1 AGNs (e.g., K.
-Inayoshi et al., 2022).
+However, one caveat in attributing the broad-line decrements to dust is the survival of dust grains in the BLRs, which likely lies within the sublimation zone (e.g., Martin1980; Kishimoto2007; Gaskell2018; Baskin2018).
+For comparison, type-1 AGNs in the SDSS typically show broad H/H ratios around 3.7, with the most extreme cases reaching 5 (Ilic2012).
+The much larger broad Balmer decrements observed in LRDs suggest that their BLRs may have extreme conditions with unusually dense gas or atypical dust properties, which differ substantially from those in typical type-1 AGNs (e.g., Inayoshi2022).
 5.2 BPT Diagram and Metallicity
-We estimate the BPT diagnostics (J.
-A.
-Baldwin et al., 1981) and the gas-phase metallicity of the narrow line regions (NLRs).
+We estimate the BPT diagnostics (Baldwin1981) and the gas-phase metallicity of the narrow line regions (NLRs).
 To measure the fluxes of forbidden lines, we fit Gaussian profiles.
@@
 In the [O I] diagram, J1025+1402 and J1022+0841 fall within the Seyfert region, while J1047+0739 lies near the edge of the H II region.
-Their positions on the [N II] and [S II] diagrams also overlap with those of low-metallicity AGNs (L.
-J.
-Kewley et al., 2013; C.
-Simmonds et al., 2016).
-Notably, the locations of these local LRDs closely match that of the cosmic-noon LRD in I.
-Juodžbalis et al.
-(2024).
+Their positions on the [N II] and [S II] diagrams also overlap with those of low-metallicity AGNs (Kewley2013; Simmonds2016).
+Notably, the locations of these local LRDs closely match that of the cosmic-noon LRD in Juodzbalis2024.
 This suggests that the narrow emission lines in both local and high-redshift LRDs are excited by similar mechanisms associated with AGN activity.
-Additionally, we detect [Fe V] 4071 emission lines in all three objects. This high-ionization line further indicates the presence of a hard ionizing radiation field and AGN activity.
 We determine the metallicities of the NLRs using the direct electron temperature () method.
@@
 The estimated (O++) is K for J1025+1402, K for J1047+0739, and K for J1022+0841.
-The of the O+ zone is estimated using the empirical (O+)–(O++) relation from Y.
-I.
-Izotov et al.
-(2006).
+The of the O+ zone is estimated using the empirical (O+)–(O++) relation from Izotov2006.
 We apply a dust attenuation correction of to J1047+0739.
-As summarized in Table 1, all three objects exhibit metal-poor NLRs with 12 + ranging from 7.42 to 8.00.
-The values are broadly consistent with those reported by Y.
-I.
-Izotov & T.
-X.
-Thuan (2008) for J1025+1402 and J1047+0739.
-We note that although the [O II] lines are detected in J1047+0739, they are significantly affected by recombination excitation enhanced by the hard ionizing radiation field (X.
-W.
-Liu et al., 2000; S.
-Tan et al., 2024).
+As summarized in Table 1, all three objects exhibit metal-poor NLRs with 12 + ranging from 7.43 to 8.01.
+The values are broadly consistent with those reported by Izotov2008 for J1025+1402 and J1047+0739.
+We note that although the [O II] lines are detected in J1047+0739, they are significantly affected by recombination excitation enhanced by the hard ionizing radiation field (Liu2000; Tan2024).
 Therefore, we do not use them for (O+) diagnostics.
-The (O+) value directly derived from the [O II] and [O II] line ratios exceeds K, approximately twice that of (O++).
-This result further suggests an AGN origin for the narrow emission lines driven by hard ionization.
+The (O+) value directly derived from the [O II] and [O II] line ratios exceeds K, more than twice the values of (O++) and the empirical (O+).
+This discrepancy suggests that [O II] is enhanced by recombination from the hard ionizing radiation field, and further supports an AGN origin for the narrow emission lines.
 Based on the Balmer decrements, BPT diagram, and direct -based metallicity, we conclude that the narrow lines in local LRDs originate from metal-poor NLRs powered by AGNs with minimal dust content.
-5.3 [Fe II] Emission Analysis
-The series of narrow [Fe II] emission lines observed in J1025+1402 and J1022+0841 stands in stark contrast to the broad Fe emission lines typically observed in type-1 AGNs (e.g., I Zw 1, M.
-P.
-Véron-Cetty et al., 2004).
-Though narrow optical [Fe II] lines have been reported in type-1 AGNs, they are blended with broad Fe lines and exhibit transitions and line ratios different from those in our sample; these narrow Fe lines are absent in type-2 AGNs (X.-B.
-Dong et al., 2010).
-To constrain the physical conditions such as temperature and density, we employ cloudy for photoionization modeling (C23, C.
-M.
-Gunasekera et al., 2023), with the latest Fe II atomic database (R.
-T.
-Smyth et al., 2019).
+5.3 [Fe II] (Fe II) Emission Analysis
+We observed a series of narrow [Fe II] and Fe II emission lines in J1025+1402 and J1022+0841 (see Figures 9 and C.1), most of which are forbidden transitions.
+These lines stand in stark contrast to the broad Fe emission lines typically observed in type-1 AGNs (e.g., I Zw 1; VMP2004).
+Though narrow optical [Fe II] (Fe II) lines have been reported in type-1 AGNs, they are blended with broad Fe lines and exhibit transitions and line ratios different from those in our sample; these narrow Fe lines are absent in type-2 AGNs (Dong2010).
+To constrain the physical conditions such as temperature and density, we employ cloudy for photoionization modeling (C23, Gunasekera2023), with the latest Fe II atomic database (Smyth2019).
 We consider scenarios in which the gas is photoionized by AGN radiation.
@@
 The AGN photoionization models yield best-fit parameters for both J1025+1402 and J1022+0841 of , , and .
-The relative [Fe II] line ratios are not strongly sensitive to metallicity, so we adopt a fiducial value of without attempting to constrain it.
-The average temperature of the [Fe II]-emitting gas is approximately K, ranging from K in the inner region to K on the surface.
-As shown in Figure 9, the best-fit model reproduces the observed [Fe II] line ratios of J1025+1402 reasonably well, although some lines (e.g., [Fe II] 5278) remain imperfectly matched.
+The relative [Fe II] (Fe II) line ratios are not strongly sensitive to metallicity, so we adopt a fiducial value of without attempting to constrain it.
+The average temperature of the [Fe II]-emitting gas777For simplicity, we hereafter refer to the gas as [Fe II]-emitting, as most emission lines are forbidden [Fe II] transitions, although a few Fe II lines are also present.
+is approximately K, ranging from K in the inner region to K on the surface.
+As shown in Figure 9, the best-fit model reproduces the observed [Fe II] (Fe II) line ratios of J1025+1402 reasonably well, although some lines (e.g., Fe II 5277) remain imperfectly matched.
 A similar level of agreement is achieved for J1022+0841 using the same best-fit parameters (Figure C.1).
 The inferred density lies between typical values for the BLRs and NLRs.
-The FWHMs of the [Fe II] lines are also intermediate ( km s-1 for J1025+1402 and km s-1 for J1022+0841), broader than the narrow H and [O III] (H –70 km s-1 in Binospec, [O III] not resolved in MODS).
-This suggests that the [Fe II] emission may arise from dense gas located outside the BLRs but inside the NLRs.
+The FWHMs of the [Fe II] (Fe II) lines are also intermediate ( km s-1 for J1025+1402 and km s-1 for J1022+0841), broader than the narrow H and [O III] (H km s-1 in Binospec, [O III] not resolved in MODS).
+This suggests that the [Fe II] (Fe II) emission may arise from dense gas located outside the BLRs but inside the NLRs.
 We emphasize that our models are qualitative.
-Fe emission lines in AGNs are inherently complex and can be influenced by a variety of physical processes, making comprehensive theoretical modeling challenging (e.g., A.
-Sarkar et al., 2021; A.
-Pandey et al., 2025).
-Further refinement is required to fully understand the origin and nature of [Fe II] emission in the LRD population.
+Fe emission lines in AGNs are inherently complex and can be influenced by a variety of physical processes, making comprehensive theoretical modeling challenging (e.g., Sarkar2021; Pandey2025).
+Further refinement is required to fully understand the origin and nature of [Fe II] (Fe II) emission in the LRD population.
 6 Case study: J1025+1402, “The Egg”
@@
 These features will be discussed in more detail below.
-An independent discovery and analysis of this object is also presented in Ji et al.
-(in prep).
+An independent discovery and analysis of this object is also presented in (Ji2025_lord).
 6.1 Strong Metal Absorption: Evidence for a Cool Gas Envelope
-The Egg exhibits exceptionally high-EW Na D 5890, 5896, K I 7665, 7699, Fe II 6148, 6244, 6456, and Ca triplets (8498, 8542, 8662; Ca T) absorption features in its optical continuum.
+6.1.1 Na D, K I, and Ca T Absorption
+The Egg exhibits exceptionally high-EW Na D 5892, 5898, K I 7667, 7701, and Ca triplets (8500, 8544, 8665; Ca T) absorption features in its optical continuum.
 We show their profiles in Figure 10 and summarize their properties in Table 4.
-Absorber | EW | FWHM |
-(Å) | (km s-1) | |
-Na D doublets | ||
-K I doublets | ||
-Fe II triplets | ||
+| Absorber | EW | FWHM | |
+| (Å) | (km s-1) | (km s-1) | |
+| Na D doublets | |||
+| K I doublets | |||
 | Ca II triplets |
-6.1.1 Na D and K I Absorption
 The Na D absorption in The Egg has an EW of Å and K I absorption shows an EW of Å.
@@
 However, the observed Na D EW in The Egg exceeds those typically found in such environments.
-The Na D EW in The Egg is higher than that observed in most galaxies and AGNs (typically Å; Y.-M.
-Chen et al., 2010; S.
-Cazzoli et al., 2016; D.
-S.
-N.
-Rupke et al., 2021; Y.
-Sun et al., 2024), and is only comparable to a few extreme IR-luminous LINERs (D.
-S.
-Rupke et al., 2005).
-In local galaxies and AGNs, high-EW Na D absorption typically traces gas outflows or inflows with velocities of several hundred km s-1 (S.
-Cazzoli et al., 2016; G.
-W.
-Roberts-Borsani & A.
-Saintonge, 2019; D.
-S.
-N.
-Rupke et al., 2021; Y.
-Sun et al., 2024), whereas in The Egg we do not observe significant velocity shifts.
-Likewise, the K I EW is orders of magnitude larger than those found in the Galactic interstellar medium or in circumstellar gas clouds around young stellar objects, where values typically range from a few up to 200 mÅ (L.
-M.
-Hobbs, 1975; D.
-E.
-Welty & L.
-M.
-Hobbs, 2001; I.
-Pascucci et al., 2015).
-All of these lines of evidence point to a non-stellar and non-ISM origin.
-6.1.2 Fe II Absorption
-The optical Fe II absorption lines at 6148, 6244, and 6456 are rarely observed in galaxies or AGNs.
-These features have been reported in a Type Ia supernova (K.
-Maguire et al., 2023), but with large blueshifts of several thousand km s-1 and broad line widths.
-In contrast, the Fe II absorption seen in The Egg shows no significant velocity shift and narrow FWHMs of 400 km s-1.
-Although the origin of these Fe II absorption features remains uncertain, their co-occurrence with Na D, K I, and Ca T absorption may suggest a shared origin.
-6.1.3 Ca T Absorption
-The Ca T absorption at 8498, 8542 and 8662 Å in The Egg has an EW of Å.
+The Na D EW in The Egg is higher than that observed in most galaxies and AGNs (typically Å; Chen2010; Cazzoli2016; Rupke2021; Sun2024), and is only comparable to a few extreme IR-luminous LINERs (Rupke2005).
+Likewise, the K I EW is orders of magnitude larger than those found in the Galactic interstellar medium or in circumstellar gas clouds around young stellar objects, where values typically range from a few up to 200 mÅ (Hobbs1975; Welty2001; Pascucci2015).
+The Ca T absorption at 8500, 8544, and 8665 Å in The Egg has an EW of Å.
 In Figure 11, we compare this value to typical Ca T EWs observed in stars, star-forming galaxies, and AGNs.
 The Ca T EW in The Egg exceeds those of most stars and all known galaxies and AGNs, where Ca T EWs are generally below 10 Å.
-Only a few individual stars, such as yellow hypergiants or supergiants (S.
-V.
-Mallik, 1997; A.
-J.
-Cenarro et al., 2001), can exhibit comparable Ca T EWs.
+Only a few individual stars, such as yellow hypergiants or supergiants (Mallik11997; Cenarro2001), can exhibit comparable Ca T EWs.
 However, reproducing the observed Ca T EWs in the integrated galaxy light would require an unusually exotic stellar population composed solely of these stars.
+All the evidence above points to an origin that is neither stellar nor associated with the ISM.
+The distinct physical characteristics of these absorbers argue against a common origin within a single cloud.
+We discuss another possible origin in §6.1.3.
+6.1.2 Other potential absorption lines
+We identify additional weak absorption features in the optical spectrum of The Egg.
+We cross-matched the visual spectral troughs against three references: the stellar absorption-line list of Struve1934, the solar spectrum line list of Moore1966, and the theoretical line catalog from the vald stellar database (Piskunov1995; Ryabchikova2015).
+As shown in Figure 12, these spectral troughs likely correspond to low-ionization absorbers of Fe I, Ti II, Mg I, Na I, Ba II, Ca I, and Ti I.
+Some of the absorption features are also reported in Ji2025_lord.
+These species all have low ionization potentials: 7.9 eV for Fe I, 13.6 eV for Ti II, 7.6 eV for Mg I, 5.1 eV for Na I, 10.0 eV for Ba II, 6.1 eV for Ca I, and 6.8 eV for Ti I.
+These absorbers are commonly observed in cool G-to-K stars (e.g., Heiter2021).
+Additionally, CH G-band absorption is present, which was also highlighted in Ji2025_lord.
+This feature is characteristic of G–K type stars and traces cool, neutral regions that favor carbon-hydrogen molecular formation.
 6.1.3 The Origin: a Cool Gas Envelope around BHs
-The unusually high-EW Na D, K I, Fe II, and Ca T absorption lines of The Egg all point to a non-stellar and non-ISM origin.
-If the optical continuum is AGN-dominated, these absorption lines likely arise from an atmosphere-like gas envelope in the vicinity of the central BHs.
-Although a stellar population struggles to explain the lines observed in The Egg, the microphysics of line formation may apply to both stellar atmospheres and LRDs.
-Therefore, a comparison to stellar spectra enables us to infer the physical conditions of the absorbing gas.
+Motivated by the unusually high-EW Na D, K I, and Ca T absorption, along with the simultaneous presence of other cool-star-like low-ionization absorptions in The Egg, we propose that these absorption lines originate from an atmosphere-like gas envelope surrounding the central BH.
+The microphysics of line formation from stellar atmospheres may likewise apply to this atmosphere-like gas envelope.
+A comparison with stellar spectra, therefore, enables us to infer the physical conditions of the absorbing gas.
 In the following, we investigate the EW of Ca T absorption as a key tracer.
 Figure 13 places The Egg’s CaT absorption in the context of stellar atmospheres.
-As shown in the left panel of Figure 12, within the stellar catalog presented here, only two metal-rich stars exhibit Ca T EWs greater than that of The Egg.
-These two stars are classified as an Ia yellow supergiant and a 0–Ia yellow hypergiant, with effective temperature () of 5500 and 5727 K, surface gravities () of 0.0 and 0.4, and metallicities [Fe/H] of and , respectively.
+As shown in the left panel of Figure 13, within the stellar catalog presented here (Mallik11997; Cenarro2001), only two metal-rich stars exhibit Ca T EWs greater than that of The Egg.
+These two stars are classified as a G0 Ia yellow supergiant and a G4 0–Ia yellow hypergiant, with effective temperature () of 5500 and 5727 K, surface gravities () of 0.0 and 0.4, and metallicities [Fe/H] of and , respectively.
 The yellow supergiant and hypergiant stars exhibit the strongest Ca T absorption due to their low surface gravity and their effective temperatures of K.
 This temperature allows Ca in the atmosphere to exist mainly in the singly-ionized state.
-In stellar atmospheres, the EWs of Ca T absorption strongly anti-correlate with surface gravity (e.g., S.
-V.
-Mallik, 1997; A.
-J.
-Cenarro et al., 2002).
+In stellar atmospheres, the EWs of Ca T absorption strongly anti-correlate with surface gravity (e.g., Mallik11997; Cenarro2002).
 This is due to the gas opacity changing with the photosphere density, which generally increases with in stellar atmosphere models.
-As decreases, the photosphere density decreases, leading to a diminished continuum absorption opacity (the main source being H- ions at and at wavelengths around – Å), while the line opacity of Ca II changes only mildly (U.
-G.
-Jorgensen et al., 1992).
+As decreases, the photosphere density decreases, leading to a diminished continuum absorption opacity (the main source being H- ions at and at wavelengths around – Å), while the line opacity of Ca II changes only mildly (Jorgensen1992).
 This enhances the contrast between the line and continuum opacities, increasing the EWs of the Ca T absorption.
@@
 If we assume a slightly lower metallicity ( to ), the inferred would be about .
-The value may be even lower given the low metallicity we have inferred from the NLR (Section 5.2).
+The value may be even lower given the low metallicity we have inferred from the NLR (§5.2).
 In conclusion, the Ca T absorption in The Egg indicates the presence of a metal-rich cool gas envelope with low surface gravity ().
@@
 The inferred value of above may not reflect the dynamical condition of the envelope, but does suggest that its atmosphere density may resemble that of very low-surface gravity stars.
-Standard stellar photosphere relations imply that at K corresponds to a photosphere density of (e.g., C.
-J.
-Hansen et al., 2004, their Equations (4.56)(4.65)).
+Standard stellar photosphere relations imply that at K corresponds to a photosphere density of (e.g., Hansen2004, their Equations (4.56)(4.65)).
 Thus, we expect the photosphere density of The Egg to be on the order of or lower than .
-We note that the FWHMs of these metal absorbers are about 300–450 km s-1 (Table 2), significantly broader than those typically observed in stellar atmospheres.
-For instance, Ca T absorption lines in stars are intrinsically narrow, with FWHMs reaching only up to 70 km s-1 in supergiant stars (A.
-V.
-Filippenko & L.
-C.
-Ho, 2003).
-The broad Ca T absorption in The Egg suggests that the cool gas envelope is turbulent or dynamically complex.
-A more detailed analysis, incorporating tailored modeling of the gas properties, is necessary for a comprehensive understanding.
-We have limited our analysis to the Ca T lines; similar atmosphere conditions are presumably responsible for the other absorption lines identified here.
+We note that these metal absorbers are offset by to km s-1 from the NLRs, with FWHMs of approximately 300–440 km s-1 (Table 4).
+It is unclear whether the offset arises from a velocity shift of the NLRs relative to the systemic redshift of the AGNs, or from the expansion of the gas envelope.
+Velocity offsets of tens to several hundred km s-1 between NLRs and AGN systemic redshifts are common in type-1 AGNs, driven by various dynamical and radiative processes (Bae2014; Shen2016).
+Meanwhile, the cool gas envelope itself is turbulent and dynamically complex.
+The FWHMs are significantly broader than those typically observed in stellar atmospheres.
+For instance, Ca T absorption lines in stars are intrinsically narrow, with FWHMs reaching only up to 70 km s-1 in supergiant stars (Filippenko2003).
+A more detailed modeling of the gas properties, dynamical state, and the stability of the envelope under the BH accretion activity is necessary for a comprehensive understanding.
+We have limited our analysis to the Ca T lines; similar atmospheric conditions are presumably responsible for the other absorption lines identified here.
 However, we do not perform the same analysis on them due to our unawareness of dedicated stellar datasets on these lines.
-Future characterization of observed stellar spectral lines combined with analysis of theoretical spectral libraries can directly test our scenario.
+Future characterization of observed stellar spectral lines, combined with analysis of theoretical spectral libraries, can directly test our scenario.
 6.2 Optical to IR Continuum: Evidence of Thermalized Emission
@@
 In this section, we investigate the potential application of atmosphere-like gas envelopes to explain the observed continuum shape of the local LRDs.
-A similar comparison conducted by A.
-de Graaff et al.
-(2025) also noted a resemblance to supergiant stars, though with notable discrepancies.
+A similar comparison conducted by deGraaff2025 also noted a resemblance to supergiant stars, though with notable discrepancies.
 In Figure 14, we compare the spectrum of The Egg to a 5000 K blackbody.
-The Å regime of The Egg’s spectrum can be reasonably well described by the Rayleigh-Jeans tail of a 5000 K blackbody.
+The Å–1.5 µm regime of The Egg’s spectrum can be reasonably well described by the Rayleigh-Jeans tail of a 5000 K blackbody.
 However, the Å region shows a significant flux deficit relative to the blue side of the blackbody spectrum, even without subtracting the extrapolation from the UV continuum as a separate component.
@@
 Such a pseudo-blackbody continuum shape signifies thermalized emission, but may involve intricate physical mechanisms.
-Recent theoretical models propose that envelope-like structures can form around SMBHs under super-Eddington accretion and emit thermal radiation with characteristic temperatures of K (D.
-Kido et al., 2025; H.
-Liu et al., 2025).
-To investigate such scenarios, we compare the observed continuum of The Egg with the models from H.
-Liu et al.
-(2025), which assume an atmosphere formed by quasi-spherical super-Eddington accretion flows.
-As shown in Figure 13, one such model from H.
-Liu et al.
-(2025) provides a good match to the optical continuum shape of The Egg.
-This model has a total luminosity and a density of at the photosphere 777 in their notation.; a star at would have this photosphere density at (C.
-J.
-Hansen et al., 2004).
+Theoretical models have proposed that envelope-like structures can form around BHs under super-Eddington accretion and emit thermal radiation (Meier1982a; Meier1982b; Zhou2019; Kido2025; Liu2025).
+To investigate such scenarios, we compare the observed continuum of The Egg with the models from Liu2025, which assume an atmosphere formed by quasi-spherical super-Eddington accretion flows.
+As shown in Figure 14, one such model from Liu2025 provides a good match to the optical continuum shape of The Egg.
+This model has a total luminosity and a density of at the photosphere 888 in their notation.; a star at would have this photosphere density at (Hansen2004).
 The photosphere (defined by the Rosseland-mean optical depth) is located at 510 AU with K.
-The frequency-dependent gas optical depth effects suppress the UV-to-optical flux, making the predicted continuum deviate from a blackbody 888We invoke no dust reddening to the model curve; we have checked that adding a modest level of dust extinction ( mag using the SMC dust extinction law) would further improve the agreement..
+The frequency-dependent gas optical depth effects suppress the UV-to-optical flux, making the predicted continuum deviate from a blackbody 999We invoke no dust reddening to the model curve; we have checked that adding a modest level of dust extinction ( mag using the SMC dust extinction law) would further improve the agreement..
 Essentially, the optical spectrum is “reddened” by the atmospheric gas; otherwise, we would have to invoke a very steep dust extinction law to explain the optical redness while retaining the declining IR slope (see Appendix D).
-However, we caution that the model assumes super-Eddington accretion, while current observations do not provide direct evidence of it.
-The integrated corresponds to about 22% of the Eddington luminosity, based on the listed in Table 1.
+However, we caution that the model of Liu2025 assumes super-Eddington accretion, while current observations do not provide direct evidence of it.
+The integrated corresponds to about 10% of the Eddington luminosity, based on the listed in Table 2.
 The true Eddington ratio is unconstrained, as both and the intrinsic bolometric luminosity are uncertain.
 In addition to the theoretical model, we present an alternative empirical parameterization of the pseudo-blackbody continuum shape in Appendix D.
-The origin of the blue UV continuum remains uncertain.
-Several observations suggest a stellar origin from young host galaxies, based on the lack of UV variability, the absence of high-ionization UV lines, and the extended, complex UV morphologies in a number of high-redshift LRDs (P.
-Rinaldi et al., 2024; H.
-B.
-Akins et al., 2025; W.
-L.
-Tee et al., 2025; R.
-P.
-Naidu et al., 2025).
-However, these features do not definitively rule out an AGN origin, either directly from the accretion disk or scattered by electrons/dust (E.
-Lambrides et al., 2024; Z.
-Li et al., 2025; K.
-Inayoshi et al., 2025).
-This is especially the case when considering the recent detection of AGN-powered high-ionization UV lines in a subset of LRDs (M.
-Tang et al., 2025).
-Recently, P.
-Rinaldi et al.
-(in prep.) found that a V-shaped continuum similar to that of LRDs originates from the nuclei of a cosmic-noon IR-bright galaxy, further suggesting that AGN emission could contribute to the blue UV continua.
-At present, we lack spectral information at rest-frame wavelengths below 2000 Å to map diagnostic UV lines that could distinguish between AGN and star-forming activity.
+The origin of the blue UV continuum ( Å) remains uncertain.
+Several observations suggest a stellar origin, rather than an AGN one, from young host galaxies, based on the lack of UV variability, the absence of high-ionization UV lines, and the extended, complex UV morphologies in a number of high-redshift LRDs (Rinaldi2024; Akins2025; Tee2025; Naidu2025; Zhuang2025).
+However, these features do not definitively rule out an AGN origin, either directly from the accretion disk or scattered by electrons/dust (Lambrides2024; Li2025; Inayoshi2025b).
+This is especially the case when considering the recent detection of AGN-powered high-ionization UV lines in a subset of LRDs (Tang2025).
+Recently, Rinaldi2025 found that a V-shaped continuum, similar to those observed in LRDs, originates from the nuclei of a cosmic-noon IR-bright galaxy.
+This finding further suggests that AGN emission could contribute to the blue UV continuum, although a circumnuclear starburst on unresolved scales could still play a role.
+At present, we lack spectral information at rest-frame wavelengths below 2000 Å to map diagnostic UV lines that could distinguish between AGN and star-forming activity (e.g., C IV, He II, N V; Feltre2016; Gutkin2016; Mingozzi2022).
 As such, the origin of this component remains inconclusive given the current data.
-Deep UV spectroscopy or polarization measurements, both of which could trace AGN light, is needed to clarify its nature.
+Deep UV spectroscopy or polarization measurements, both of which could trace AGN light, are needed to clarify its nature.
 6.3 The Dust Emission
 The WISE photometry of local LRDs suggests that the dust content in these objects, while not as strong as implied by a standard AGN dust torus, is also not negligible.
-A typical AGN dust torus, when scaled to LRD luminosities, would show rapidly increasing emission beyond 1–2 µm, peaking around 10 µm, which is clearly not the case for LRDs (M.
-Stalevski et al., 2016; B.
-Wang et al., 2025b).
+A typical AGN dust torus, when scaled to LRD luminosities, would show rapidly increasing emission beyond 1–2 µm, peaking around 10 µm, which is clearly not the case for LRDs (Stalevski2016; BWang2025).
 All three local LRDs show flattening or an increase in flux beyond rest-frame 10 µm (W3 and W4 bands; Figure 1).
-At , this wavelength range is beyond JWST coverage, and it is either only partially covered by Spitzer and Herschel or falls into gaps between their coverage, leaving it largely unconstrained for high-redshift LRDs (H.
-B.
-Akins et al., 2024; C.
-M.
-Casey et al., 2024, 2025; D.
-J.
-Setton et al., 2025).
+At , this wavelength range is beyond JWST coverage, and it is either only partially covered by Spitzer and Herschel or falls into gaps between their coverage, leaving it largely unconstrained for high-redshift LRDs (Akins2024; Casey2024; Casey2025; Setton2025).
 Qualitatively, the IR excess from the WISE photometry in these local LRDs (Figure 1) reveals at least dust components with temperatures approximately between and K.
 The emission at rest-frame 3 µm corresponds to the peak of a K blackbody, and the emission at 10 µm matches a K blackbody.
-As shown in Figure 13, if assuming templates from J.
-Lyu & G.
-H.
-Rieke (2021) that include dust components at , 1038, 300, and 90 K (see Appendix §D.2), the IR emission of The Egg can be well reproduced by significant dust components at 1038 K and 300 K.
+As shown in Figure 14, if assuming templates from Lyu2021 that include dust components at , 1038, 300, and 90 K (see Appendix §D.2), the IR emission of The Egg can be well reproduced by significant dust components at 1038 K and 300 K.
 However, the derived dust properties depend sensitively on the chosen templates.
-We discuss alternative dust templates based on Haro 11 (J.
-Lyu et al., 2016) and the degeneracies in Appendix D.3.
-In the absence of longer-wavelength data, it is difficult to robustly constrain the dust properties, particularly the cold dust content with temperatures below 100 K.
+We discuss alternative dust templates based on Haro 11 (Lyu2016) and the degeneracies in Appendix D.3.
+In the absence of continuous near- to mid-infrared coverage and far-infrared constraints, it is challenging to robustly constrain the dust structures and grain properties in these systems.
+Recent IR and sub-mm observations have placed first constraints on the dust content of high-redshift LRDs (Casey2024; Casey2025; Setton2025; Xiao2025; Li2025; Chen2025).
+The weak IR emission observed in the three local LRDs in this work is consistent with both the detections and upper limits reported for high-redshift LRDs.
+Deep ALMA non-detections and stacking analyses indicate modest dust masses () and peak temperatures of 100–300 K (Casey2025; Setton2025) for high-redshift LRDs.
+These inferred temperatures agree with the dust components detected in the three local LRDs at rest-frame wavelengths µm.
+For dust attenuation, Chen2025 placed stringent upper limits on high-redshift LRDs (), whereas other studies have inferred non-negligible values () from broad-line Balmer decrements or continuum fitting (Ji2025; D'Eugenio2025_z7p04; Taylor2025), underscoring a tension in current estimates.
+We do not attempt to precisely quantify the attenuation on the continuum or broad lines in this work.
+Proper modeling of dust attenuation in such atypical environments requires careful treatment of multiple factors, including the geometry and distribution of dust needed to maintain nearly dust-free NLRs, uncertainties in broad-line Balmer decrements, the survivability of dust grains in BLRs, and the validity of energy balance along the line of sight.
+We also emphasize the need for longer-wavelength follow-up observations of local LRDs, such as with ALMA.
+These observations can yield more stringent constraints on the dust content than those currently available for their high-redshift counterparts, pending the arrival of next-generation IR facilities (e.g., PRIMA, Moullet2023).
 7 A conceptual picture of LRDs
 | Physical Process | Observational Evidence | Caveat and prospects |
-BH accretion | Broad emission lines (e.g., Balmer and Paschen series, infrared O I, infrared Fe II) with FWHM 1000 km s-1 | Optically thick gas envelope prevents direct observations of BH accretion process.
+| BH accretion | Broad emission lines (e.g., Balmer and Paschen series, infrared O I, infrared Fe II) with FWHM 1000 km s-1 | Optically thick gas envelope prevents direct observations of BH accretion process (e.g., emission directly from the accretion disks).
 |
-AGN Emission Line Regions | Narrow and broad line regions consistent with AGN ionization; [Fe II] emission lines with FWHM and inferred density between broad and narrow components | Modeling of [Fe II] is complex and potentially degenerate |
-Cool gas envelope | Metal absorption lines, especially with extremely high-EW Na D, K I, Fe II, and Ca T absorption features in The Egg; the thermalized emission in good agreement with theoretical atmosphere models (H.
-Liu et al., 2025) | Requires detailed modeling of opacity, cooling, and circum-BH environments; must also account for the FWHMs of absorption lines |
+| AGN Emission Line Regions | Narrow and broad line regions consistent with AGN ionization; [Fe II] and Fe II emission lines with FWHM and inferred density between broad and narrow components | Modeling of [Fe II] and Fe II is complex and potentially degenerate |
+| Cool gas envelope | Cool-star like metal absorption features, especially extremely high-EW Na D, K I, and Ca T absorption lines in The Egg; the thermalized emission in good agreement with theoretical atmosphere models (Liu2025) | Envelopes are dynamically complex, requiring models that capture absorption-line FWHMs and velocities, as well as detailed prescriptions of opacity, cooling, and circum-BH environments.
+|
 | Non-spherical geometry and viewing angle | Optically thin channels within the envelope are required for AGN ionizing photons to escape and illuminate the BLRs, intermediate [Fe II] regions, and NLRs; may also explain the diversity in continuum shapes and the presence or absence of high-EW metal absorbers.
@@
 |
-Outflowing optically-thick gas | Blueshifted Balmer and He I absorption lines | Their correlation with other properties of LRDs (i.e., Balmer break strength) remains uncertain; the origin of He I emission (ionized gas versus dense neutral gas) requires further study.
+| Outflowing/inflowing optically-thick gas | Blueshifted Balmer and He I absorption lines | Their correlation with other properties of LRDs (i.e., Balmer break strength, continuum shape) remains uncertain; the origin of He I emission (ionized gas versus dense neutral gas) requires further study.
 |
-Dust emission | near-IR excess in WISE photometry (rest-frame 3-20 µm) | Constraints on dust components depend on model assumptions; cannot fully characterize dust properties from the AGN to the ISM due to the lack of longer-wavelength data.
+| Dust emission | near- to mid-IR excess in WISE photometry (rest-frame 3-20 µm) | Constraints on dust components depend on model assumptions; cannot fully characterize dust properties from the AGN to the ISM due to the lack of longer-wavelength data.
 |
@@
 BLRs and NLRs.
-The broad Balmer and Paschen lines, along with the infrared O I emission lines (in The Egg, see Figure 2), indicate that the observed line broadening arises from gas motions within the BLRs surrounding the BHs.
+The broad Balmer and Paschen lines, along with the infrared O I µm emission lines (in The Egg, see Figure 2), indicate that the observed line broadening arises from gas motions within the BLRs surrounding the BHs.
 Analysis of the Balmer decrements, BPT diagrams, and metallicity suggests that the narrow emission lines originate from AGN-powered NLRs with low metallicity and minimal dust.
-[Fe II] emitting gas intermediate between the BLRs and NLRs.
-The AGN-excited [Fe II] emission lines indicate gas with a density of approximately cm-3, intermediate between the BLRs and NLRs.
-The FWHM of the [Fe II] lines ( 200 km s-1 for The Egg and 300 km s-1 for J1022+0841 in MODS spectra) is likewise between that of the broad ( km s-1) and narrow lines (17 km s-1 for The Egg and 77 km s-1 for J1022+0841 in the Binospec spectra).
+The high decrements of broad Balmer lines (H/H) imply either high-density gas with complex radiative transfer effects or atypical dust properties in BLRs.
+[Fe II]-emitting gas intermediate between the BLRs and NLRs.
+The AGN-excited [Fe II] (or Fe II) emission lines indicate gas with a density of approximately cm-3, intermediate between the BLRs and NLRs.
+The FWHM of the [Fe II] lines ( 100-300 km s-1) is likewise between that of the broad ( km s-1) and narrow lines ( km s-1).
+Determining whether these regions can emit other lines requires deeper, higher-resolution spectroscopy to accurately decompose broad lines, such as those in the Balmer and Paschen series, or to detect fainter features.
 The absence of these [Fe II] lines in J1047+0739 may be attributed to a lower covering fraction of the intermediate emitting regions.
 A cool gas envelope with 5000–6000 K and photosphere density of ().
-The optical–IR continua of the three local LRDs exhibit pseudo-blackbody shapes indicative of thermalized emission and are in good agreement with theoretical atmosphere models (H.
-Liu et al., 2025).
-The presence of extremely high-EW metal absorbers (Na D, K I, Fe II, Ca T) in The Egg further indicates a cool, optically thick gas envelope surrounding the BH.
+The optical–IR continua of the three local LRDs exhibit pseudo-blackbody shapes indicative of thermalized emission and are in good agreement with theoretical atmosphere models (Liu2025).
+The presence of extremely high-EW metal absorbers (Na D, K I, Ca T) in The Egg, along with other potential G-to-K star-like low-ionization absorbers (Figure 12), further indicates a cool, optically thick gas envelope surrounding the BH.
 Analysis of the Ca T EW, by analogy with stellar atmospheres, suggests an atmospheric temperature of approximately 5000–6000 K and surface gravity of at the photosphere, corresponding to a density of .
@@
 A consequence of an optically thick cool gas envelope is that the BH accretion process, i.e., the emission from the accretion disk itself, is not directly observable.
+The origin of the UV emission remains inconclusive with the current data, and we do not attempt to address it within this framework.
+It could originate either from stars in the host galaxy or from AGN light scattered by the surrounding gas.
+We leave a more thorough investigation of its nature to future work and observations.
 Non-spherical geometry and role of the viewing angle.
@@
 However, we emphasize that this is a speculative scenario, and a larger statistical sample is needed to better understand the roles of geometry and viewing angle.
-Differences in the metallicities of the cool gas envelop would also contribute to the diversity of observed metal absorption features.
-Optically thick outflows.
-In most high-redshift LRDs, the Balmer absorption lines are blueshifted, with only a few cases showing redshifted H absorption with minimal velocity offsets (X.
-Lin et al., 2024; X.
-Ji et al., 2025; F.
-D’Eugenio et al., 2025).
-Up to now, all detected He I absorbers in LRDs are blueshifted (I.
-Juodžbalis et al., 2024; B.
-Wang et al., 2025a; F.
-Loiacono et al., 2025).
+Differences in the metallicities of the cool gas envelope would also contribute to the diversity of observed metal absorption features.
+Optically thick outflows/inflows.
+In most high-redshift LRDs, the Balmer absorption lines are blueshifted, with only a few cases showing redshifted H absorption with minimal velocity offsets (Lin2024; Ji2025; DEugenio2025).
+Up to now, all detected He I absorbers in LRDs are blueshifted (Juodzbalis2024; Wang2025; Loiacono2025).
 The Balmer absorbers are superimposed on broad Balmer emission lines, and the He I absorbers are superimposed on both the broad emission and the underlying continuum.
 We thus attribute these absorbers to optically thick outflowing gas located either within or outside the BLRs.
-The rare instances of redshifted Balmer absorption with small velocity shifts may indicate gas cooling and inflowing back toward the nucleus.
-Although the relationship between the absorption and LRD properties (e.g., Balmer break strength) remains unclear, it is a plausible connection that merits further exploration.
+Balmer absorption arises from high-density neutral gas.
+He I absorbers may originate from either high-density neutral gas clouds responsible for the Balmer absorption or partially ionized gas clouds with lower density.
+The rare instances of redshifted Balmer absorption with small velocity shifts, like those seen in The Egg and J1022+0841, may indicate inflows, which could represent accreting material or gas cooling and moving back toward the nucleus.
+The relationship between these absorptions and other properties of LRDs (e.g., Balmer break strength, continuum shape) is not yet established.
+However, since their presence can be attributed to high-density hydrogen ( cm-3; Inayoshi2024), there may be plausible connections that merit further investigation.
 Dust at – K.
 The WISE photometry of the three local LRDs is consistent with dust components with temperatures ranging from to K.
-These dust components, cooler than typical AGN-heated hot dust but warmer than ISM dust, is likely located outside the 5000 K atmosphere yet still within the AGN vicinity.
+These dust components, cooler than typical AGN-heated hot dust but warmer than ISM dust, are likely located outside the 5000 K atmosphere yet still within the AGN vicinity.
 However, we note that the analysis of the dust components is highly uncertain, model-dependent, and subject to degeneracies.
+Due to the lack of observational constraints, the exact dust geometry and grain properties remain unknown.
+Future modeling efforts with improved IR observations are needed.
 In the proposed scenario, the BH resides in an environment distinct from that of typical type-1 AGNs.
@@
 For the three objects, if we take the total integrated luminosity , assuming that the UV continuum is AGN light, the Eddington ratio would be .
-The current observations do not provide sufficient evidence to determine whether LRDs are in a super-Eddington accretion mode, as proposed by some theoretical scenarios (e.g., K.
-Inayoshi et al., 2024; F.
-Pacucci & R.
-Narayan, 2024).
-The accretion rate of the central BH and its disk can significantly influence the geometry of both the disk and the surrounding gas envelope (e.g., C.
-Dotan & N.
-J.
-Shaviv, 2011).
+The current observations do not provide sufficient evidence to determine whether LRDs are in a super-Eddington accretion mode, as proposed by some theoretical scenarios (e.g., Inayoshi2024a; Pacucci2024; Madau2025).
+The accretion rate of the central BH and its disk can significantly influence the geometry of both the disk and the surrounding gas envelope (e.g., Dotan2011).
 We leave a detailed exploration of this to future work.
@@
 We searched the SDSS spectroscopic database for local LRDs using high-redshift LRD templates and identified three LRDs at –.
+Their discovery suggests a conservative lower limit for LRD number density at of .
 We followed up these sources with MMT/Binospec, LBT/MODS, and Magellan/FIRE spectroscopy, which reveal detailed spectral features that probe the physical nature of these local LRDs.
@@
 The three local objects closely resemble the high-redshift LRDs.
-They display compact morphologies.
-HST narrow-band imaging of J1047+0739 reveals that both the continuum and the H line emission are AGN-dominated, with the compact yet clumpy host galaxy (half-light radius about 530 pc) contributing a minor fraction to the continuum.
-All the three objects exhibit clear V-shaped UV-to-optical continua with inflections around the Balmer break, and a decreasing IR continuum from Å to 2 µm.
+They display compact morphologies (0.85-1.6 pkpc).
+HST narrow-band imaging of J1047+0739 reveals that both the continuum and the H line emission are AGN-dominated, with the compact yet clumpy host galaxy ( pc) contributing a minor fraction to the continuum.
+All three objects exhibit clear V-shaped UV-to-optical continua with inflections near the Balmer break, and a decreasing IR continuum from Å to 2 µm.
 All of them show broad H emission with FWHMs km s-1, along with broad Paschen and He I lines.
-All of them exhibit blueshifted He I 10830 absorbers; two of them show H absorbers, and one even shows a redshifted H absorber without an H counterpart.
+All of them exhibit blueshifted He I 10833 absorbers; two of them show H absorbers, and one even shows a redshifted H absorber without an H counterpart.
 None of the objects shows significant optical or IR variability.
 Two were targeted in hard X-rays but were either not detected or detected with low statistical significance.
-The Balmer and Paschen decrements, covering from H to H12 and from Pa to Pa12 (or Pa9 for one object), indicate negligible dust attenuation in the NLRs.
-The broad H/H ratios are in all three objects, higher than those observed in most local AGNs, indicating either unusually high-density gas or atypical dust properties in the BLRs.
-The BPT diagram shows that the narrow emission lines are photoionized by AGNs and overlap with the region occupied by a known LRD. The NLRs are metal-poor, with 12+(O/H)7.43–8.00.
-Two of the objects exhibit prominent narrow [Fe II] emission lines that are distinct from those typically seen in type-1 AGNs.
+The Balmer and Paschen decrements, covering from H to H12 and from Pa to Pa12 (or Pa9 for one object), indicate negligible dust attenuation in the NLRs, consistent with Case B recombination.
+The broad H/H ratios are 8-34 in the three objects, higher than those observed in most local AGNs, indicating either unusually high-density gas or atypical dust properties in the BLRs.
+The BPT diagram shows that the narrow emission lines are photoionized by AGNs, with the three LRDs occupying positions consistent with those reported by Juodzbalis2024.
+Their NLRs are metal-poor, with 12 + (O/H) 7.43–8.01.
+Two of the objects exhibit prominent narrow [Fe II] (or Fe II) emission lines that are distinct from those typically seen in type-1 AGNs.
 Photoionization modeling suggests that the [Fe II]-emitting gas is AGN-powered, with physical conditions characterized by , , , and an average temperature of K.
 Based on the inferred gas density and FWHMs, we propose that the [Fe II]-emitting region lies at an intermediate scale between the BLRs and NLRs.
-One of the local LRDs, J1025+1402 (The Egg) exhibits exceptionally high EWs in Na D, K I, Fe II, and Ca T absorption lines, with narrow FWHMs of about 300–400 km s-1 and no significant velocity shifts.
+One of the local LRDs, J1025+1402 (The Egg) exhibits exceptionally high EWs in Na D, K I, and Ca T absorption lines, with FWHMs of about 300–440 km s-1 and no significant velocity shifts.
 Such strong metal absorbers are rarely seen in the ISM, galaxies, AGNs, or stars.
 Its Ca T EW is matched only by that of two supergiant and hypergiant stars.
+The Egg also exhibits potential low-ionization absorbers, such as Fe I, Ti II, Mg I, Ba II, etc., characteristic of G- to K-type stars.
 These extreme properties suggest a non-stellar and non-ISM origin.
-We interpret the Ca T absorption in The Egg as arising from a cool gas envelope around the BH, analogous to stellar atmospheres.
-We infer – and at the photosphere, corresponding to a density of g cm-3.
+We interpret these absorption features in The Egg as arising from a cool gas envelope around the BH, analogous to stellar atmospheres.
+Based on its Ca T EW, we infer – and at the photosphere, corresponding to a density of g cm-3.
 The optical-IR continua of these local LRDs exhibit pseudo-blackbody shapes.
-The shapes are consistent with theoretical models featuring atmospheres surrounding BHs (H.
-Liu et al., 2025), with characteristic temperatures of K.
+The shapes are consistent with theoretical models featuring atmospheres surrounding BHs (Liu2025), with characteristic temperatures of K.
 The WISE photometry indicates modest dust emission at temperatures of – K.
@@
 The overall AGN structure consists of BLRs, an intermediate region that contains [Fe II]-emitting gas, and NLRs with minimal dust attenuation.
-The geometry of these structures, along with the viewing angle, may account for the diverse observed features across different objects.
+The geometry of these structures, along with the viewing angle, may account for the diverse observed features across different objects (i.e., the presence or absence of cool star–like absorption).
 Local LRDs offer a valuable opportunity to investigate the physical nature of this population.
@@
 Do LRDs, and the scenario illustrated in Figure 15, represent a specific evolutionary phase of SMBHs?
+What is the accretion mode of these objects (super-Eddington or sub-Eddington)?
 Are they connected to conventional type-1 and type-2 AGNs?
@@
 While answering these questions is beyond the scope of this paper, they are critical for understanding the broader context of BHs and merit future investigation.
-Ongoing wide-field surveys, such as DESI and Euclid, along with upcoming missions like the Roman Space Telescope, will open new windows for identifying LRDs at low redshift.
+Ongoing wide-field surveys, such as DESI and Euclid, along with upcoming missions such as the Prime Focus Spectrograph (PFS), LSST, and Roman Space Telescope, will open new windows for identifying LRDs at low redshift.
 These efforts will enable studies of their cosmic evolution and contributions to the history of SMBH growth.
@@
 It indicates that the H line is primarily contributed by the AGN.
-We model the morphology of the continuum map using galfit (C.
-Y.
-Peng et al., 2002).
-We assume a PSF and four Sérsic components.
+We model the morphology of the continuum map using galfit (Peng2002).
+We assume a PSF and four Sérsic components, with the number of Sérsic components determined through several trials; four Sérsic yielded the optimal fit.
 The right panel of Figure A.1 presents the best-fit model and the PSF-subtracted components, which represent the host galaxy.
-The host galaxy has compact, yet clumpy morphology, with a half-light radius of 530 pc.
+The host galaxy has compact, yet clumpy morphology, with a half-light radius of 568 pc.
 The luminosity of the central PSF is 5.5 times that of the host galaxy, indicating that the AGN contributes about 85% of the total continuum light in the optical.
@@
 Appendix B Emission line fitting
-For strong emission lines with visible broad components (e.g., H, Pa), we model each line with a single narrow Gaussian plus one or more broad Gaussian or Lorentzian profiles to capture the composite broad features.
+For the H emission lines of The Egg and J1022+0841, we use observations from LBT/MODS and MMT/Binospec.
+Given the higher resolution of Binospec (–5300, compared to –1300 for MODS), we adopt the Binospec spectra for the fiducial fits and degrade the best-fit models to the MODS resolution for consistency checks.
+In the Binospec spectra, the absorption and emission components are fitted simultaneously, with a linear polynomial included to account for the local continuum.
+The [N II] 6550, 6585 doublet is modeled jointly with H using two narrow Gaussians, with their FWHMs tied to that of the narrow H component and their flux ratio fixed to 3.
+As shown in Figure B.1, the best-fit models reproduce the Binospec spectra well.
+When degraded to the resolution of MODS, the models remain in good agreement with the MODS H profiles.
+Small discrepancies are present in the broad wings of The Egg, arising from the fact that the instrumental line-spread functions might not be perfectly Gaussian.
+We verify that such discrepancies introduce only a 5% difference in the broad H luminosity when fitting the MODS spectrum alone, and thus have a negligible impact on the derived physical properties (e.g., ).
+For J1022+0841, we test two different fitting models: one assuming a single Gaussian absorber superimposed on the broad emission lines, and the other using two Gaussian absorbers.
+As shown in Figure B.1, the two-Gaussian model provides a better fit, more accurately capturing the wing of the absorption profile.
+The fitting results using the two-Gaussian model are reported in Table 3.
+For the single-Gaussian model, we obtain km s-1, FWHM= km s-1, and EW= Å.
+The normalized H absorption profiles shown in Figure 6 are derived by dividing the absorbed broad H emission by the intrinsic broad H line profile.
+For strong emission lines exhibiting visible broad components (e.g., H, Pa) in the LBT/MODS and Magellan/FIRE spectra, each line is modeled with a single narrow Gaussian plus one or more broad Gaussian or Lorentzian profiles to capture the composite broad features.
 A linear polynomial is added to account for the local continuum.
-Specifically, broad Paschen lines are modeled with Lorentzian profiles, leveraging the high-resolution FIRE spectra to resolve line wings, while broad H components are depicted by two to three Gaussian profiles.
-The [N II] 6548, 6583 doublet is fit simultaneously with H.
-We model them with two narrow Gaussians, fixing their FWHMs to match the narrow H component and setting their flux ratio to 3.
+Specifically, broad Paschen lines are modeled with Lorentzian profiles, leveraging the high-resolution FIRE spectra to resolve line wings.
 In J1047+0739, the Paschen lines and adjacent IR He I lines are modeled simultaneously, with the He I emission represented by broad and narrow Gaussian components.
 Note that for each line profile fit, the parameters of the narrow Gaussian component are treated independently, as the spectral resolution varies across wavelengths and instruments.
-For example, [O III]5007 in the MODS spectra is unresolved, with its observed FWHM determined by the wavelength-dependent line spread function.
+For example, [O III] 5008 in the MODS spectra is unresolved, with its observed FWHM determined by the wavelength-dependent line spread function.
 Therefore, we do not use its fit to constrain the narrow components of H or Pa.
-For the H emission lines of The Egg and J1022+0841, whether from MODS or Binospec spectra, we simultaneously fit the absorption and emission profiles.
-Given the higher resolution of the Binospec spectra, we rely on them for more precise absorber measurements.
-As shown in Figure B.1, a Gaussian absorption component (EW Å) is applied on top of the composite broad H profile (i.e., the sum of the two broad Gaussians).
-For J1022+0841, we test two different fitting models: one assuming a single Gaussian absorber superimposed on the broad emission lines, and the other using two Gaussian absorbers.
-As shown in Figure B.1, the two-Gaussian model provides a better fit, more accurately capturing the profile wing.
-Since the He I 10830 absorption profiles are complex, we model the intrinsic emission by masking the absorption-affected regions.
-The remaining line profile wings, unaffected by absorption, are used to constrain and recover the symmetric emission line profiles.
+The measurements of key emission lines are summarized in Table B.1.
+| J1025+1402 | J1047+0739 | J1022+0841 | |
+| ( erg s-1) | |||
+| (km s-1) | – | – | |
+| ( erg s-1) | |||
+| ( erg s-1) | |||
+| ( erg s-1) | |||
+| ( erg s-1) | |||
+| FWHMHβ,broad (km s-1) | |||
+| ( erg s-1) | |||
+| FWHMPaα,broad (km s-1) | |||
+| ( erg s-1) | |||
+| FWHMPaα,narrow (km s-1) | |||
+| ( erg s-1) | |||
+| FWHMPaβ,broad (km s-1) | a | ||
+| ( erg s-1) | |||
+| FWHMPaβ,narrow (km s-1) | – a | ||
+| ( erg s-1) | |||
+| FWHMPaγ,broad (km s-1) | |||
+| ( erg s-1) | |||
+| FWHMPaγ,narrow (km s-1) | |||
+| ( erg s-1) | |||
+| FWHMPaδ,broad (km s-1) | |||
+| ( erg s-1) | |||
+| FWHMPaδ,narrow (km s-1) | – |
+For H line of The Egg, we mask the absorbed region and fit the intrinsic emission profile.
+Similarly, the intrinsic He I 10833 emission is modeled by excluding regions affected by absorption.
+The unaffected line wings are used to constrain and recover the symmetric emission profile.
 The emission is modeled with one or two broad Gaussian components plus a narrow component.
-For H line of The Egg, we also mask the absorbed region and fit the intrinsic emission profile.
-The normalized H absorption profiles shown in Figure 6 are derived by dividing the absorbed broad H emission by the intrinsic broad H line profile.
-For H in The Egg and He I 10830 in all three objects, the profiles are normalized by the total unabsorbed profiles, including both the broad emission lines and continua.
-Normalizing H using only the broad emission line is tested but results in an absorption peak below zero.
-We then model the normalized H absorber with a single Gaussian and the normalized He I 10830 absorber with multiple Gaussians.
+In Figure 6, the H profile in The Egg and the He I 10833 profiles in all three objects are normalized by the total unabsorbed profiles, including both the broad emission lines and continua.
+Normalizing H using only the broad emission line is tested, but results in an absorption peak below zero.
+We then model the normalized H absorber with a single Gaussian and the normalized He I 10833 absorber with multiple Gaussians.
 Table 3 summarizes the best-fit parameters for each absorber.
-Name | H | H | He I 10830 | ||||||
-FWHM | EW | FWHM | EW | FWHM | EW | ||||
-(km s-1) | (km s-1) | (Å) | (km s-1) | (km s-1) | (Å) | (km s-1) | (km s-1) | (Å) | |
-The Egg | 73 | ||||||||
-(J1025+1402) | |||||||||
-total | |||||||||
-J1047+0739 | – | – | – | – | – | – | |||
-total | |||||||||
-J1022+0841 | – | – | – | (two Gaussian) | |||||
-total | |||||||||
-(one Gaussian) | |||||||||
 Appendix C [Fe II] modeling
-To model AGN photoionization scenarios for [Fe II] emission, we adopt AGN SEDs from C.
-Jin et al.
-(2012).
+To model AGN photoionization scenarios for [Fe II] emission, we adopt AGN SEDs from Jin2012.
 While the chosen SED corresponds to an Eddington ratio of 0.17, our cloudy modeling is intended to be qualitative and is not highly sensitive to this assumed value.
@@
 To determine the best-fit models, we compare the observed emission line fluxes with those predicted by the modeled [Fe II] spectrum.
-Specifically, we use the fluxes of [Fe II] 4235, 4244, 4278, 4287, 4416, 5158, 5262, 7155, and 7172.
-Additionally, since [Fe II] 5169 is highly sensitive to the gas conditions but is not detected, we include its 3 upper limit as an extra constraint.
-All line fluxes, both observed and modeled, are normalized to that of [Fe II] 4244, and we compute the value for each model grid.
+Specifically, we use the fluxes of Fe II 4234, [Fe II] 4245, 4278, 4289, 4416 (blend of 4415 and 4418), 5159, 5263, 7157, and 7174.
+Additionally, since Fe II 5171 is highly sensitive to the gas conditions but is not detected, we include its 3 upper limit as an extra constraint.
+All line fluxes, both observed and modeled, are normalized to that of [Fe II] 4245, and we compute the value for each model grid.
 For comparison, we also explored models using a 5000 K blackbody as the ionizing source.
 The temperature is chosen based on our analysis in §6 and §7.
-These models fail to reproduce the observed [Fe II] line ratios, with significant discrepancies particularly in [Fe II] 4235, 4244, 4278, 4287, and 4416.
+These models fail to reproduce the observed [Fe II] line ratios, with significant discrepancies particularly in Fe II 4234, [Fe II]4245, 4278, 4289, and 4416.
 In Figure C.1, we present the best-fit [Fe II] model for J1022+0841. Its best-fit parameters are identical to those used for J1025+1402.
@@
 D.1 Theoretical atmosphere model + dust templates
-We begin with the best-matched theoretical models from H.
-Liu et al.
-(2025) and model the dust emission using the dust templates from J.
-Lyu & G.
-H.
-Rieke (2021).
-These dust templates, originally developed for type-1 AGNs, include four components: blackbodies at 2100 K, 1038 K, and 300 K, and a graybody at 90 K.
+We begin with the best-matched theoretical models from Liu2025 and model the dust emission using the dust templates from Lyu2021.
+As demonstrated in Lyu2017 and Lyu2021, typical AGN-heated dust emission SED can be well can be well characterized by the combination of four major dust components: blackbodies at 2100 K, 1038 K, and 300 K, and a graybody at 90 K.
 We extrapolate the theoretical model spectra by assuming a blackbody tail in the near-IR, consistent with the model’s effective temperature.
@@
 Figure 14 demonstrates that this model successfully reproduces the near-infrared photometry of The Egg.
-Figure D.1 shows the SEDs of J1047+0739 and J1022+0841 compared with and the models from H.
-Liu et al.
-(2025) that visually match the data.
-The intrinsic luminosity of the H.
-Liu et al.
-(2025) model is erg s-1 for J1047+0739, with a reference density of g cm-3 (which is comparable to the photosphere density; see H.
-Liu et al.
-(2025) for exact definition).
+Figure D.1 shows the SEDs of J1047+0739 and J1022+0841 compared with and the models from Liu2025 that visually match the data.
+The intrinsic luminosity of the Liu2025 model is erg s-1 for J1047+0739, with a reference density of g cm-3 (which is comparable to the photosphere density; see Liu2025 for exact definition).
 These model parameters give an effective temperature of K.
 For J1022+0841, the model luminosity is erg s-1 with g cm-3, and the resulting K.
-We note that the SED of J1047+0739 is less well reproduced, mainly due to the non-negligible galaxy light.
-The HST narrow-band imaging (see Appendix A) reveals that, although the AGN dominates the continuum, the host galaxy contributes at least 18% of the continuum light around the H wavelengths.
+We note that the SED of J1047+0739 is less well reproduced, possibly due to the non-negligible galaxy light.
+Indeed, the HST narrow-band imaging (see Appendix A) reveals that, although the AGN dominates the continuum, the host galaxy contributes about 15% of the continuum light around the H wavelengths.
 The contribution from the host galaxy is likely minor in The Egg and J1022+0841, both of which exhibit more prominent V-shaped SEDs with redder optical continua.
@@
 This parameterization is purely descriptive, intended to capture the observed continuum shape but not to provide a physical interpretation.
-We adopt the functional form of a modified Calzetti law (S.
-Noll et al., 2009), but it should not be interpreted as indicating any particular amount of dust attenuation.
+We adopt the functional form of a modified Calzetti law (Noll2009), but it should not be interpreted as indicating any particular amount of dust attenuation.
 The parameterization can be expressed as:
@@
 It represents a distinct component, which could be either due to the AGN light or the host galaxy.
-denotes the form of the modified Calzetti law (S.
-Noll et al., 2009), where represents a generalized attenuation curve (not necessarily tied to physical dust properties), and is the power-law modification parameter, normalized at 5500Å.
+denotes the form of the modified Calzetti law (Noll2009), where represents a generalized attenuation curve (not necessarily tied to physical dust properties), and is the power-law modification parameter, normalized at 5500 Å.
 For J1047+0739, we introduce an additional free parameter to model the Balmer jump as a discontinuity in the power-law at 3645 Å.
 is the blackbody spectrum at temperature , and is its amplitude.
-For the dust emission, we adopt the dust templates from J.
-Lyu & G.
-H.
-Rieke (2021).
+For the dust emission, we adopt the dust templates from Lyu2021.
 and are the amplitude and blackbody or gray body spectrum of the -th dust component.
 The best-fit continua of the three local LRDs are shown in Figure D.2.
-The corresponding best-fit parameters and component luminosities are summarized in Table 5.
+The corresponding best-fit parameters and component luminosities are summarized in Table D.1.
 We do not include associated uncertainties given the descriptive nature of this parameterization.
@@
 In J1047+0739, the W2 photometry at a rest-frame wavelength of 4 µm is not well reproduced, as simple blackbody or graybody models fail to capture features such as PAH emission and silicate absorption.
-In J1022+0841, a dust component with K appears, but this primarily reflects the influence of the four dust templates adopted from J.
-Lyu & G.
-H.
-Rieke (2021).
-To explore alternative interpretations, we replace the dust component in Equation D1 with the Haro 11 dust template from J.
-Lyu et al.
-(2016) and fit the continua of the three local LRDs, as shown in Figure D.3.
+In J1022+0841, a dust component with K appears, but this primarily reflects the influence of the four dust templates adopted from Lyu2021.
+To explore alternative interpretations, we replace the dust component in Equation D1 with the Haro 11 dust template from Lyu2016 and fit the continua of the three local LRDs, as shown in Figure D.3.
 Notably, the dust emission in J1022+0841 is well reproduced by the Haro 11 template, suggesting that its dust may reside in the ISM and be heated by AGN-driven feedback or massive star clusters in the host galaxy.